### v4

# Computation method for perfectoid purity and perfectoid BCM-regularity

Abstract.
In this paper, we introduce a criterion for a sufficient condition of perfectoid purity and a computation method for perfectoid pure threshold of ${\rm div}(p)$.
1. Introduction
Recently, the theory of singularities in mixed characteristic has been developed in many papers [MSTWW], [BMPSTWW24], [HLS]. The theory has applications for both the theory of algebraic geometry and the ring theory in mixed characteristic, for example, the minimal model theory [BMPSTWW23], [TY23], and the Skoda-type theorem [HLS]. In this paper, we study the class of singularities perfectoid purity [p-pure] and perfectoid BCM-regularity [bcm-reg], which are known as analogues of singularities in positive characteristic $F$-purity and strongly $F$-regular, respectively. It is known that for a Noetherian local ring $(R,\mathfrak{m})$ with $p\in\mathfrak{m}$, if $R$ is complete intersection and $R/pR$ is $F$-pure or strongly $F$-regular, then $R$ is perfectoid pure or perfectoid BCM-regular, respectively. On the other hand, we have a useful criterion called “Fedder’s criterion” [Fedder], [Glassbrenner] for $F$-purity and strongly $F$-regularity of complete intersection local rings, thus many examples can be constructed combining them.
In this paper, we construct new examples of perfectoid pure or perfectoid BCM-regular rings. Let $(R,\mathfrak{m})$ be a $p$-torsion free Noetherian local ring such that $p\in\mathfrak{m}$ and $\overline{R}:=R/pR$ is $F$-finite. We say that $R$ is perfectoid pure if there exists a pure extension $R\to B$ to a perfectoid. If $R$ is perfectoid pure, the perfectoid pure threshold $\operatorname{ppt}(R;{\rm div}(p))$ of ${\rm div}(p)$ is defined by
$\inf\left\{i/p^{e}\geq 0\middle|\begin{array}[]{l}\textup{there exist a perfectoid $R$-algebra $B$ and $\varpi$ a $p^{e}$-power}\\ \textup{root of $p$ such that $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure}\end{array}\right\},$
see Proposition 2.9 for details. It is known [p-pure]*Theorem 6.6 that if $R$ is complete intersection and $\overline{R}$ is $F$-pure, then $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1$ (cf. Remark 2.10). The first purpose of the paper is to construct examples of perfectoid pure rings $R$ with $\operatorname{ppt}(R;{\rm div}(p))\neq 1$. For this purpose, we define the ring homomorphism $\Phi_{R,n}\colon\overline{R}\to\overline{W}_{n}(R):=W_{n}(R)/pW_{n}(R)$, where $W_{n}(R)$ is the ring of Witt vectors over $R$ of length $n$, by
$\Phi_{R,n}\colon\overline{R}\to\overline{W}_{n}(R)\ ;\ a\mapsto(\widetilde{a}^{p},0,\ldots,0),$
where $\widetilde{a}\in R$ is a lift of $a$ in $R$ (see Proposition 4.1 for well-definedness). We regard $\overline{W}_{n}(R)$ as an $\overline{R}$-module via $\Phi_{R,n}$ and denote it by $Q_{R,n}$, then we note that $Q_{R,1}=F_{*}\overline{R}$. We introduce the following notions.
Definition 1.1.
(Definition 4.4, cf. [Yobuko19])
  • Let $n$ be a positive integer. We say that $R$ is $n$-quasi-$F$-split if $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$ splits as $\overline{R}$-modules and $R$ is quasi-$F$-split if $R$ is $m$-quasi-$F$-split for some positive integer $m$.
  • The quasi-$F$-splitting height $\mathrm{ht}(R)$ is defined by
[…]
    if $R$ is quasi-$F$-split, and $\mathrm{ht}(R):=\infty$ if $R$ is not quasi-$F$-split.
  • Let $n,e$ be positive integers. We define $\overline{R}$-module $Q_{R,n,e}$ by the following pushout diagram as $\overline{R}$-modules;
[…]
    and the induced $\overline{R}$-module homomorphism is denoted by
    $\Phi_{R,n,e}\colon R\xrightarrow{\Phi_{R,n}}Q_{R,n}\to Q_{R,n,e}.$
  • We say that $R$ is quasi-$(F,F^{\infty})$-split if $R$ is quasi-$F$-split and $\Phi_{R,\mathrm{ht}(R),e}$ splits as $\overline{R}$-modules for every positive integer $e\geq 1$.
The first main result in this paper is Theorem A.
Theorem A.
We assume $R$ is complete intersection. Let $n$ be a positive integer.
  1. (1)
    $\operatorname{ht}(R)=n$ if and only if $R$ is perfectoid pure and
    $1-\frac{1}{p}-\cdots-\frac{1}{p^{n-1}}\geq\operatorname{ppt}(R;{\rm div}(p))\geq 1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}$
    (Theorem 5.15). In particular, if $R/pR$ is quasi-$F$-split, then $R$ is perfectoid pure (Theorem 5.12).
  2. (2)
    $R$ is quasi-$(F,F^{\infty})$-split of height $n$ if and only if $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1-1/p-\cdots-1/p^{n-1}$ (Corollary 5.16).
Furthermore, we obtain a Fedder type criterion for quasi-$F$-splitting (Theorem 4.12) and obtain examples as follows (cf. Example 5.9).
  1. (1)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z]]/(z^{2}+y^{3}+z^{5})$, then
    $\operatorname{ppt}(R;{\rm div}(p))=\begin{cases}1/8&\textup{if $p=2$},\\ 5/9&\textup{if $p=3$},\\ 4/5&\textup{if $p=5$}.\end{cases}$
  2. (2)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z,w]]/(w^{2}+xyz(x+y+z))$ and $p=2$, then $\operatorname{ppt}(R;{\rm div}(p))=1/2$. We note that $\overline{R}$ is not quasi-$F$-split by [kty]*Example 7.11.
  3. (3)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z,w]]/(w^{2}+xyz(x+y+z)+p(xy+xz+yz)w)$ and $p=2$, then $\operatorname{ppt}(R;{\rm div}(p))=1/4$.
  4. (4)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z]]/(z^{2}+x^{2}y+xy^{n})$ for $n\geq 2$ and $p=2$, then $\operatorname{ppt}(R;{\rm div}(p))=1/2^{\lceil\log_{2}n\rceil}$.
The next main result Theorem B is on the perfectoid BCM-regularity ([bcm-reg]*Definition 6.9) for graded rings.
Theorem B.
(Theorem 6.3) Let $S=\bigoplus_{m\geq 0}S_{m}$ be a Noetherian graded ring such that $S$ is $p$-torsion free and $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $S_{0}$. We assume $\overline{S}:=S/pS$ is normal quasi-Gorenstein and
$a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{\dim{\overline{S}}}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}<0.$
We further assume that $\operatorname{Spec}{\overline{S}}\backslash\{\overline{S}_{>0}\}$ is strongly $F$-regular. If $S_{(p,S_{>0})}$ is quasi-$F$-split, then $S_{(p,S_{>0})}$ is perfectoid BCM-regular. Furthermore, the converse implication holds if $p=2$.
Moreover, we obtain a Fedder type criterion for quasi-$F$-splitting (Theorem 4.12), thus Theorem 4.12 is a criterion for perfectoid BCM-regularity for rings satisfying the assumption in Theorem B in characteristic two. By the inversion of adjunction type result (Proposition 4.6), if $\overline{S}$ is quasi-$F$-split, so is $S$. Furthermore, we construct examples as follows.
Theorem C.
(Corollary 6.6, cf. Corollaries 6.4 and 6.5) There exists a Noetherian graded ring $S=\bigoplus_{i\in\mathbb{Z}_{\geq 0}}S_{i}$ such that $S$ satisfies the following properties.
  1. (1)
    $S$ is torsion free and $S_{0}$ is a divisorially valuation ring with maximal ideal $(p)$.
  2. (2)
    We set $\mathfrak{m}:=(S_{>0},p)$ then $S_{\mathfrak{m}}$ is perfectoid BCM-regular.
  3. (3)
    $S_{\mathfrak{m}}/pS_{\mathfrak{m}}$ is normal Gorenstein, but not strongly $F$-regular.
Finally, we compute a perfectoid pure threshold for quasi-$F$-split graded rings with $a(\overline{S})=0$.
Theorem D.
(Theorem 6.9) Let $S=\bigoplus_{m\geq 0}S_{m}$ be a Noetherian graded ring such that $S$ is $p$-torsion free and $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $S_{0}$. We assume $S_{\mathfrak{m}}$ is complete intersection and
$a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{\dim{\overline{S}}}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}=0,$
where $\mathfrak{m}:=(S_{>0},p)$. If $\operatorname{ht}(\overline{S_{\mathfrak{m}}})=n<\infty$, then $S_{\mathfrak{m}}$ is perfectoid pure and we have
$\operatorname{ppt}(S_{\mathfrak{m}};{\rm div}(p))=1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}.$
Furthermore, we obtain an interesting example of non-perfectoid pure ring as follows.
Example 1.2.
(Proposition 6.10) Let $R^{\prime}:=\mathbb{Z}_{p}[[x,y,z]]/(x^{3}+y^{3}+z^{3})$ and $p\equiv 2\ \textrm{mod}\ 3$. Then $R^{\prime}$ is perfectoid pure but $R^{\prime}\otimes_{\mathbb{Z}_{p}}R^{\prime}$ is not perfectoid pure.
A key ingredient of all main results is running a program (Definition 5.4) to obtain a sequence of elements of $H^{d}_{\mathfrak{m}}(\overline{R})$. Whether or not this program stops is closely related to the singularity of a given ring (cf. Remark 5.5(3) and Proposition 5.7).
Acknowledgements.
The author wishes to express his gratitude to Ryo Ishizuka, Teppei Takamastu, Kenta Sato, and Tatsuro Kawakami for valuable discussion. He is also grateful to Linquan Ma for helpful comments. The author was supported by JSPS KAKENHI Grant number JP20J11886.
2. Preliminary
2.1. Notation
In this subsection, we summarize the notation and terminologies used in this paper.
  1. (1)
    Throughout this paper, we fix a prime number $p$ and we set $\mathbb{F}_{p}:=\mathbb{Z}/p\mathbb{Z}$.
  2. (2)
    For a ring $R$ of characteristic $p>0$, $F:R\to R$ denotes the absolute Frobenius ring homomorphism, that is, $F(r)=r^{p}$ for every $r\in R$.
  3. (3)
    For a ring $R$ of characteristic $p>0$, we say that $R$ is $F$-finite if $F$ is finite.
  4. (4)
    For a ring $R$ of characteristic $p>0$, we set $R_{perf}:=\mathrm{colim}_{F}R$ and call it by the perfection of $R$.
  5. (5)
    Throughout this paper, we use the terminology perfectoid by means of [BMS19]*Definition 3.5. For a semiperfectoid $R$, the perfectoidazation is denoted by $R_{perfd}$ (cf. [BS]*Section 7).
  6. (6)
    Let $R$ be a ring and $f$ an element of $R$. A sequence $\{f^{1/p^{e}}\}_{e\geq 0}$ is called a compatible system of $p$-power roots of $f$ if $(f^{1/p^{e+1}})^{p}=f^{1/p^{e}}$ for every $e\geq 0$ and $f^{1/p^{0}}=f$. Furthermore, the ideal $(f^{1/p^{e}}\mid e\geq 0)$ of $R$ is denoted by $(f^{1/p^{\infty}})$.
  7. (7)
    For a ring $R$ and a ring homomorphism $\phi\colon R\to R$, we say that $\phi$ is a lift of Frobenius if $\phi(a)\equiv a^{p}\ \textrm{mod}\ p$ for every $a\in R$.
  8. (8)
    For a ring homomorphism $f\colon A\to B$ and a finitely generated ideal $I$ of $A$, we say that $f$ is $I$-completely faithfully flat if for every $I$-torsion $A$-module $N$, we have $\operatorname{Tor}_{i}^{A}(B,N)=0$ for every $i\neq 0$ and $B/IB$ is faithfully flat over $A/I$. If $I=(p)$, we call it by $p$-completely faithfully flat.
  9. (9)
    For a ring homomorphism $f\colon A\to B$, we say that $f$ is pure if for every $A$-module $M$, the homomorphism $M\to M\otimes_{A}B$ is injective.
  10. (10)
    Let $(R,\mathfrak{m})$ be a local ring and $B$ an $R$-algebra. We say that $B$ is big Cohen-Macaulay (BCM) algebra over $R$ if for every system of parameter $x_{1},\ldots,x_{d}$ of $R$, the sequence $x_{1},\ldots,x_{d}$ is a regular sequence on $B$. Furthermore, we say that $R$ is perfectoid BCM-regular if every extension $R\to B^{\prime}$ to a perfectoid BCM $R$-algebra $B^{\prime}$ is pure ([bcm-reg]*Definition 6.9).
2.2. The ring of Witt vectors
We recall the definition of the ring of Witt vectors. The basic reference is [Serre]*II, § 6.
Lemma 2.1.
For $n\in\mathbb{Z}_{\geq 0}$, we define the polynomial $\varphi_{n}\in\mathbb{Z}[X_{0},\ldots X_{n}]$ by
$\varphi_{n}:=\sum_{i=0}^{i=n}p^{i}X_{i}^{p^{n-i}}.$
Then, there exist polynomials $S_{n},P_{n}\in\mathbb{Z}[X_{0},\ldots,X_{n},Y_{0},\ldots,Y_{n}]$ for any $n\in\mathbb{Z}_{\geq 0}$ such that
$\varphi_{m}(S_{0},\ldots,S_{m})=\varphi_{m}(X_{0},\ldots X_{m})+\varphi_{m}(Y_{0},\ldots Y_{m})$
and
$\varphi_{m}(P_{0},\ldots,P_{m})=\varphi_{m}(X_{0},\ldots X_{m})\cdot\varphi_{m}(Y_{0},\ldots Y_{m})$
hold for any $m\in\mathbb{Z}_{\geq 0}$.
Proof.
See [Serre]*II, Theorem 6. ∎
Definition 2.2.
Let $A$ be a ring.
  • We set
    $W(A):=\{(a_{0},\ldots,a_{n},\ldots)|a_{n}\in A\}=\prod_{\mathbb{Z}_{\geq 0}}A,$
    equipped with the addition
    $(a_{0},\ldots,a_{n},\ldots)+(b_{0},\ldots,b_{n},\ldots):=(S_{0}(a_{0},b_{0}),\ldots,S_{n}(a_{0},\ldots,a_{n},b_{0},\ldots,b_{n}),\ldots)$
    and the multiplication
    $(a_{0},\ldots,a_{n},\ldots)\cdot(b_{0},\ldots,b_{n},\ldots):=(P_{0}(a_{0},b_{0}),\ldots,P_{n}(a_{0},\ldots,a_{n},b_{0},\ldots,b_{n}),\ldots).$
    Then $W(A)$ is a ring with $1=(1,0,\ldots)\in W(A)$, which is called the ring of Witt vectors over $A$. For $a\in A$, we set $[a]:=(a,0,\ldots)\in W(A)$. We define an additive map $V$ by
    $V\colon W(A)\rightarrow W(A);(a_{0},a_{1},\ldots)\mapsto(0,a_{0},a_{1},\ldots).$
  • We define the ring $W_{n}(A)$ of Witt vectors over $A$ of length $n$ by
    $W_{n}(A):=W(A)/V^{n}W(A).$
    We note that $V^{n}W(A)\subseteq W(A)$ is an ideal. For $a\in A$ and $n\in\mathbb{Z}_{\geq 1}$, we set $[a]:=(a,0,\ldots,0)\in W_{n}(A)$. We define a $W(A)$-module homomorphism $R$ by
    $W(A)\rightarrow W_{n}(A);(a_{0},a_{1},\ldots)\mapsto(a_{0},a_{1},\ldots,a_{n-1})$
    and call $R$ the restriction map. Moreover, we define
    $V\colon W_{n}(A)\rightarrow W_{n+1}(A),$
    and
    $R\colon W_{n+1}(A)\rightarrow W_{n}(A),$
    in a similar way. The image of $\alpha\in W(A)$ or $W_{n}(A)$ by $V$ is denoted by $V\alpha$.
  • For any ring homomorphism $f\colon A\to A^{\prime}$, a ring homomorphism
    $W(A)\to W(A^{\prime})\ ;\ (a_{0},a_{1},\ldots)\mapsto(f(a_{0}),f(a_{1}),\ldots)$
    is induced and the induced homomorphism is commutative with $R$ and $V$. Note that $f$ also induces a ring homomorphism $W_{n}(A)\to W_{n}(A^{\prime})$.
  • For an ideal $I$ of $A$, we denote the kernels of $W_{n}(A)\to W_{n}(A/I)$ and $W_{n}(A)\to W_{n}(A/I)$ induced by the natural quotient map $A\to A/I$ are denoted by $W(I)$ and $W_{n}(I)$, respectively.
  • We define a ring homomorphism $\varphi$ by
    $\varphi\colon W(A)\rightarrow\prod_{\mathbb{Z}_{\geq 0}}A;(a_{0},\ldots,a_{n},\ldots)\mapsto(\varphi_{0}(a_{0}),\ldots,\varphi_{n}(a_{0},\ldots,a_{n}),\ldots),$
    where $\prod_{\mathbb{Z}_{\geq 0}}A$ is the product in the category of rings. The map $\varphi$ is injective if $p\in A$ is a non-zero divisor. The element $\varphi_{n}(a_{0},\ldots,a_{n})$ (resp. the element $\varphi(a_{0},\ldots a_{n},\ldots)$) is called the ghost component (resp. the vector of ghost components) of $(a_{0},\ldots a_{n},\ldots)\in W(A)$.
Proposition 2.3.
Let $R$ be a ring. Then we have the following properties:
  1. (1)
    For $a,b\in R$ $m\in\mathbb{Z}_{\geq 0}$, we have $[a]V^{m}([b])=V^{m}([a^{p^{m}}b])$ in $W(R)$
  2. (2)
    For $\alpha,\beta\in W(R)$, we have $V(\alpha)V(\beta)=pV(\alpha\beta)$ in $W(R)$.
Proof.
We take a ring homomorphism $\pi\colon A:=\mathbb{Z}[X_{f}\mid f\in R]\to R$ such that $\pi(X_{f})=f$, then $\pi$ is surjective. Since $\pi$ induces a ring homomorphism $W(\pi)\colon W(A)\to W(R)$ and $A$ is $p$-torsion free, we may assume that $R$ is $p$-torsion free. We take $\varphi\colon W(R)\to\prod_{\mathbb{Z}_{\geq 0}}R$ as in Definition 2.2, then $\varphi$ is injective. Let $s\colon\prod_{\mathbb{Z}_{\geq 0}}R\to\prod_{\mathbb{Z}_{\geq 0}}R$ be a map defined by $s(a_{0},a_{1},\ldots)=(0,a_{0},a_{1},\ldots)$, then we have $s(\alpha\beta)=s(\alpha)s(\beta)$ for $\alpha,\beta\in\prod_{\mathbb{Z}_{\geq 0}}R$. Let us show (1). We have
$\varphi([a]V^{m}([b]))$ $=\varphi([a])\varphi(V^{m}([b]))=(a,a^{p},\cdots)p^{m}s^{m}(b,b^{p},\ldots)$
$=p^{m}s^{m}(a^{p^{m}}b,a^{p^{m+1}}b^{p},\ldots)=\varphi(V^{m}([a^{p^{m}}b])),$
for $a,b\in R$. Since $\varphi$ is injective, we obtain the assertion (1). Next, let us show (2). Furthermore, we have $\varphi\circ V(\alpha)=ps\circ\varphi(\alpha)$ for $\alpha\in W(R)$, thus
$\varphi(V(\alpha)V(\beta))=\varphi\circ V(\alpha)\varphi\circ V(\beta)=p^{2}s\circ\varphi(\alpha)s\circ\varphi(\beta)=p^{2}s\circ\varphi(\alpha\beta)=pV(\alpha\beta)$
for $\alpha,\beta\in W(R)$, as desired. ∎
2.3. Test perfectoid and perfectoid pure threshold
In this subsection, we introduce the notion of test perfectoid and perfectoid pure threshold of ${\rm div}(p)$. The existence of the test perfectoids is proved in Appendix A.
Definition 2.4.
(cf. Theorem A.1) Let $R$ be a ring and $R_{\infty}$ a perfectoid $R$-algebra. We say that $R_{\infty}$ is a test perfectoid over $R$ if for every perfectoid $R$-algebra $B$, there exist a $p$-completely faithfully flat $R$-algebra homomorphism $B\to B^{\prime}$ to a perfectoid $R$-algebra $B^{\prime}$ and an $R$-algebra homomorphism $R_{\infty}\to B^{\prime}$.
Remark 2.5.
Let $R\to S$ be a ring homomorphism of perfectoids. Then $R\to S$ is $p$-completely faithfully flat if and only if $R/p^{n}\to S/p^{n}$ is faithfully flat for each positive integer $n$ by [BS]*Lemma 3.8(2) and [BMS19]*Corollary 4.8(3).
Proposition 2.6.
Let $R$ be a ring, and $R_{\infty}$ a perfectoid over $R$.
  1. (1)
    Let $R_{\infty}\to R^{\prime}_{\infty}$ be a $p$-completely faithfully flat $R$-algebra homomorphism to a perfectoid $R^{\prime}_{\infty}$. Then $R_{\infty}^{\prime}$ is a test perfectoid over $R$.
  2. (2)
    Let $I$ be an ideal of $R$. Then $(R_{\infty}/IR_{\infty})_{perfd}$ is a test perfectoid over $R/I$.
Proof.
Let us show (1). We take a perfectoid $R$-algebra $B$, then replacing $B$ by a $p$-completely faithfully flat extension of $B$, we may assume $B$ is $R_{\infty}$-algebra. By [CS]*Proposition 2.1.11(b), the $p$-adic completion $V^{\prime}:=B\widehat{\otimes}_{R_{\infty}}R^{\prime}_{\infty}$ of $B\otimes_{R_{\infty}}R^{\prime}_{\infty}$ is a perfectoid. Since $R_{\infty}\to R^{\prime}_{\infty}$ is $p$-completely faithfully flat, so is $B\to B^{\prime}$ by Remark 2.5. Since there exists an $R$-algebra homomorphism $R^{\prime}_{\infty}\to B^{\prime}$, as desired. Next, let us show (2). We take a perfectoid $R/I$-algebra $B$. Since $R_{\infty}$ is a test perfectoid, there exists a $p$-completely faithfully flat extension $B\to B^{\prime}$ of perfectoid $R$-algebras such that there exists an $R$-algebra homomorphism $R_{\infty}\to B^{\prime}$. Since $B^{\prime}$ is an $R/I$-algebra and $R_{\infty}$-algebra, it is an $R_{\infty}/IR_{\infty}$-algebra. Thus, we obtain the $R/I$-algebra homomorphism $(R_{\infty}/IR_{\infty})_{perfd}\to B^{\prime}$, as desired. ∎
Lemma 2.7.
Let $R$ be an $\mathbb{F}_{p}$-algebra, $R_{\infty}$ a test perfectoid over $R$. Let $M$ be an $R$-module. Then we have
$\mathrm{Ker}(M\to R_{\infty}\otimes_{R}M)=\mathrm{Ker}(M\to R_{perf}\otimes_{R}M),$
where $R_{perf}$ is the perfection of $R$.
Proof.
Since $R_{\infty}$ is a perfectoid $\mathbb{F}_{p}$-algebra, it is a perfect ring. Therefore, there exists an $R$-algebra homomorphism $R_{\mathrm{perf}}\to R_{\infty}$, thus we have
$\mathrm{Ker}(M\to R_{perf}\otimes_{R}M)\subseteq\mathrm{Ker}(M\to R_{\infty}\otimes_{R}M).$
On the other hand, since $R_{\infty}$ is a test perfectoid, there exist a faithfully flat extension $R_{\mathrm{perf}}\to B$ to a perfect $R$-algebra and an $R$-algebra homomorphism $R_{\infty}\to B$. Therefore we have
$\mathrm{Ker}(M\to R_{\infty}\otimes_{R}M)\subseteq\mathrm{Ker}(M\to B\otimes_{R}M)\overset{(*)}{=}\mathrm{Ker}(M\to R_{perf}\otimes_{R}M),$
where ()(*)( ∗ ) follows from the faithfully flatness of $R_{\mathrm{perf}}\to B$. ∎
Definition 2.8.
Let $R$ be a $p$-torsion free Noetherian ring with $p$ in its Jacobson radical. We assume $R$ is perfectoid pure. We define the perfectoid pure threshold $\operatorname{ppt}(R;{\rm div}(p))$ of ${\rm div}(p)$ by
$\inf\left\{i/p^{e}\geq 0\middle|\begin{array}[]{l}\textup{there exist a perfectoid $R$-algebra $B$ and $\varpi$ a $p^{e}$-power}\\ \textup{root of $p$ such that $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure}\end{array}\right\}.$
Proposition 2.9.
Let $R$ be a $p$-torsion free Noetherian ring with $p$ in its Jacobson radical. Then we obtain the following assertions.
  1. (1)
    The ring $R$ is perfectoid pure if and only if every test perfectoid $R_{\infty}$ over $R$ is pure.
  2. (2)
    Let $R_{\infty}$ be a test perfectoid over $R$ and $\{p^{1/p^{e}}\}$ a compatible system of $p$-power roots of $p$ in $R_{\infty}$. If $R$ is perfectoid pure and $R_{\infty}$ is $p$-torsion free, then we have
    $\operatorname{ppt}(R;{\rm div}(p))=\inf\{\alpha\in\mathbb{Z}[1/p]_{\geq 0}\mid\textup{$R\to R_{\infty}\xrightarrow{\cdot p^{\alpha}}R_{\infty}$ is pure }\}.$
  3. (3)
    We assume $(R,\mathfrak{m})$ is a complete local ring. We take $S\to R$ and $R^{S_{\infty}}_{perfd}$ be as in [p-pure]*Lemma 4.23. Then we have
    $\operatorname{ppt}(R;{\rm div}(p))=\inf\{\alpha\in\mathbb{Z}[1/p]_{\geq 0}\mid\textup{$R\to R^{S_{\infty}}_{perfd}\xrightarrow{\cdot p^{\alpha}}R^{S_{\infty}}_{perfd}$ is pure }\}.$
Proof.
We assume $R$ is perfectoid pure and take a test perfectoid $R\to R_{\infty}$ over $R$. Since $R$ is perfectoid pure, there exists a pure extension $R\to B$ such that $B$ is a perfectoid. Since $R_{\infty}$ is a test perfectoid, we may assume that there exists an $R$-algebra homomorphism $R_{\infty}\to B$ by the proof of [p-pure]*Lemma 4.5. Therefore, $R\to R_{\infty}$ is also pure. On the other hand, by Theorem A.1, $R$ has a test perfectoid, thus we obtain the converse implication.
Next, let us show assertion (2). By definition, we have
$\inf\{\alpha\in\mathbb{Z}[1/p]_{\geq 0}\mid\textup{$R\to R_{\infty}\xrightarrow{\cdot p^{\alpha}}R_{\infty}$ is pure }\}\geq\operatorname{ppt}(R;{\rm div}(p)).$
We take $i,e\in\mathbb{Z}_{\geq 1}$, a perfectoid $R$-algebra $B$, and $p^{e}$-th root $\varpi$ of $p$ such that $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure. It is enough to show that $R\to R_{\infty}\xrightarrow{\cdot p^{i/p^{e}}}R_{\infty}$ is pure. By the proof of (1) and the definition of test perfectoid, we may assume there exists an $R$-algebra homomorphism $\nu\colon R_{\infty}\to B$. By [BMS18]*Lemma 3.10, we have $\varpi B=\nu(p^{1/p^{e}})B$, thus there exists a unit $u\in B$ such that $\varpi=u\nu(p^{1/p^{e}})$, and in particular, we have $R\to B\xrightarrow{\cdot\nu(p^{1/p^{e}})}B$ is pure. By the commutative diagram
[…]
we obtain that $R\to R_{\infty}\xrightarrow{\cdot p^{1/p^{e}}}R_{\infty}$ is pure.
Finally, assertion (3) follows from the same argument as in the proof of assertion (2) and the proof of [p-pure]*Lemma 4.23. ∎
Remark 2.10.
Let $(R,\mathfrak{m})$ be a $p$-torsion free Noetherian local ring with $p\in\mathfrak{m}$. Then $\operatorname{ppt}(R;{\rm div}(p))=1$ if and only if $(R,p)$ is perfectoid pure by means of [p-pure]*Definition 6.1. Indeed, we take a test perfectoid $R_{\infty}$ over $R$ such that $R_{\infty}$ contains a compatible system $\{p^{1/p^{e}}\}$ of $p$-power root of $p$ and $R_{\infty}$ is $p$-torsion free, where the existence follows from Theorem A.1. Since $R$ and $R_{\infty}$ are $p$-torsion free, the pair $(R,p)$ is perfectoid pure if and only if $R\to R_{\infty}\xrightarrow{\cdot p^{1-1/p^{e}}}R_{\infty}$ is pure for every $e\geq 1$, thus it is equivalent to $\operatorname{ppt}(R;{\rm div}(p))=1$ by Proposition 2.9. In particular, if $R/pR$ is $F$-pure and $R$ is complete intersection, then $\operatorname{ppt}(R;{\rm div}(p))=1$ by [p-pure]*Theorem 6.6.
3. Witt vectors of rings with Frobenius lifts
In this section, we study the structure of the ring of Witt vectors of rings with Frobenius lift. Theorem 3.7 is a key ingredient of results in the paper.
Notation 3.1.
Let $A$ be a $p$-torsion free ring with a ring homomorphism $\phi\colon A\to A$ such that $\phi(a)\equiv a^{p}\ \textrm{mod}\ p$ for every $a\in A$. The ring homomorphism
$\varphi:=\prod\varphi_{r}\colon W(A)\to\prod_{\mathbb{N}}A$
is defined as in Definition 2.2.
Proposition 3.2.
(cf. [Ill79]*Section 1.3) We use Notation 3.1. There exists a ring homomorphism
$s_{\phi}\colon A\to W(A)$
such that
$\varphi_{n}(s_{\phi}(a))=\phi^{n}(a).$
The composition $A\xrightarrow{s_{\phi}}W(A)\to W_{n}(A)$ is also denoted by $s_{\phi}$.
Proof.
It is enough to show that for every $a\in A$, there exist $a_{0},a_{1},\ldots$ such that
$\sum_{0\leq s\leq n}p^{s}a_{s}^{p^{n-s}}=\phi^{n}(a).$
Then if we define $s_{\phi}(a)=(a_{0},a_{1},\ldots)$, we have $\varphi_{n}(s_{\phi}(a))=\phi^{n}(a)$. Since $\varphi$ is injective, $s_{\phi}$ is a ring homomorphism.
Therefore, we prove the existence of $a_{0},a_{1},\ldots$ as above by the induction on $n$. First, we define $a_{0}:=a$. Next, we assume $n\geq 1$ and there exist $a_{0},a_{1},\ldots,a_{n-1}$ such that
$\sum_{0\leq s\leq n-1}p^{s}a_{s}^{p^{n-s-1}}=\phi^{n-1}(a).$
Then we have
$\phi^{n}(a)=\sum_{0\leq s\leq n-1}p^{s}\phi(a_{s})^{p^{n-s-1}}$
Since $\phi(a_{s})\equiv a_{s}^{p}\ \textrm{mod}\ p$, we have $p^{s}\phi(a_{s})^{p^{n-s-1}}\equiv p^{s}a_{s}^{p^{n-s}}\ \textrm{mod}\ p^{n}$. Therefore, we have
$\phi^{n}(a)\equiv\sum_{0\leq s\leq n-1}p^{s}a_{s}^{p^{n-s}}\ \textrm{mod}\ p^{n},$
and there exists $a_{n}$ such that
$\phi^{n}(a)=\sum_{0\leq s\leq n}p^{s}a_{s}^{p^{n-s}},$
as desired. ∎
Proposition 3.3.
We use Notation 3.1. We regard $W_{n}(A)$ as $A$-module by $s_{\phi}$ for every positive integer $n$. Then $V\colon\phi_{*}W_{n-1}(A)\to W_{n}(A)$ is an $A$-module homomorphism for $n\geq 2$.
Proof.
First, we note that the diagram
${\phi_{*}W_{n-1}(A)}$${W_{n}(A)}$${\phi_{*}A\oplus\cdots\oplus\phi^{n}_{*}A}$${A\oplus\cdots\oplus\phi^{n}_{*}A}$$\scriptstyle{V}$$\scriptstyle{\phi_{*}\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{V^{\prime}}$
commutes, where $V^{\prime}$ is defined by $V^{\prime}((a_{0},\ldots,a_{n-1}))=(0,pa_{0},\ldots,pa_{n-1})$. Furthermore, since $V^{\prime}$ is an $A$-module homomorphism and the $A$-module homomorphisms $\phi_{*}\varphi$, $\varphi$ are injective, the map $V$ is $A$-module homomorphism, as desired. ∎
Definition 3.4.
We use Notation 3.1. We define a map
$\Delta_{W_{n}}\colon W_{n}(A)\to W_{n-1}(A)$
by
$V\Delta_{W_{n}}((a_{0},\ldots,a_{n-1}))=(a_{0},\ldots,a_{n-1})-s_{\phi}(a_{0}).$
Since $s_{\phi}(a_{0})-[a_{0}]$ is contained in the image of $V$, the map $\Delta_{W_{n}}$ is well-defined.
Proposition 3.5.
We use Notation 3.1. Maps $\Delta_{W_{n}}$ are group homomorphisms and satisfy the following properties
  1. (1)
    $\Delta_{W_{n}}(s_{\phi}(a))=0$ for $a\in A$ ,
  2. (2)
    $\Delta_{W_{n}}(V\alpha)=\alpha$ for $\alpha\in W_{n-1}(A)$, and
  3. (3)
    $\Delta_{W_{n}}(\alpha\beta)\equiv s_{\phi}\circ\phi(b)\Delta_{W_{n}}(\alpha)+s_{\phi}\circ\phi(a)\Delta_{W_{n}}(\beta)\ \textrm{mod}\ pW_{n-1}(A)$ for $\alpha,\beta\in W_{n}(A)$ and $a:=R^{n-1}(\alpha),b:=R^{n-1}(\beta)$.
Proof.
We take an element $a\in A$, then
$V\Delta_{W_{n}}(s_{\phi}(a))=s_{\phi}(a)-s_{\phi}(a)=0,$
thus we have $\Delta_{W_{n}}(s_{\phi}(a))=0$. The second assertion is clear by definition of $\Delta_{W_{n}}$. For the third condition, we consider
$V\Delta_{W_{n}}(\alpha\beta)$ $=\alpha\beta-s_{\phi}(ab)$
$=(\alpha-s_{\phi}(a))\beta+s_{\phi}(a)(\beta-s_{\phi}(b))$
$=s_{\phi}(b)V\Delta_{W_{n}}(\alpha)+s_{\phi}(a)V\Delta_{W_{n}}(\beta)+(V\Delta_{W_{n}}(\alpha))(V\Delta_{W_{n}}(\beta))$
=$\overset{(\star)}{=}V(s_{\phi}\circ\phi(b)\Delta_{W_{n}}(\alpha)+s_{\phi}\circ\phi(a)\Delta_{W_{n}}(\beta)+p\Delta_{W_{n}}(\alpha)\Delta_{W_{n}}(\beta)),$
where ()($\star$)( ⋆ ) follows from Proposition 3.3 and Proposition 2.3 (2). Therefore, we have
$\Delta_{W_{n}}(\alpha\beta)=s_{\phi}\circ\phi(b)\Delta_{W_{n}}(\alpha)+s_{\phi}\circ\phi(a)\Delta_{W_{n}}(\beta)+p\Delta_{W_{n}}(\alpha)\Delta_{W_{n}}(\beta),$
as desired. ∎
Theorem 3.6.
(c.f. [kty]*Lemma 3.17) We use Notation 3.1. Let $a\in A$. We define the map
$\Delta_{s}(a)=\Delta_{W_{2}}\circ\Delta_{W_{3}}\circ\cdots\circ\Delta_{W_{s+1}}([a])\in A$
for $s\geq 1$, and $\Delta_{0}(a)=a$. Then for every $a\in A$, we have
$[a]=s_{\phi}(a)+Vs_{\phi}(\Delta_{1}(a))+\cdots+V^{n-1}s_{\phi}(\Delta_{n-1}(a))+V^{n}\Delta_{n}(a).$
In particular, we have the following two equations
$a^{p^{n}}$ $=\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)$
$\Delta_{n}(a)$ $=\frac{\Delta_{1}(a^{p^{n-1}})}{p^{n-1}}.$
Furthermore, we obtain
$\Delta_{n}(a)\equiv\begin{cases}a^{p^{n}-p}\Delta_{1}(a)+a^{p^{n}-2p}\Delta_{1}(a)^{p}\ \textrm{mod}\ p&p=2\\ a^{p^{n}-p}\Delta_{1}(a)\ \textrm{mod}\ p&\textup{otherwise}\end{cases}$
for integers $n\geq 2$.
Proof.
By the definition of $\Delta_{W_{n+1}}([a])$, we have
$[a]=s_{\phi}(a)+V\Delta_{W_{n+1}}([a]).$
Furthermore, since the first component of $\Delta_{W_{n+1}}([a])$ is $\Delta_{1}(a)$, we have
$\Delta_{W_{n+1}}([a])=s_{\phi}(\Delta_{1}(a))+V\Delta_{W_{n}}(\Delta_{W_{n+1}}([a])).$
We note that the first component of $\Delta_{n-r}\circ\cdots\Delta_{n+1}([a])$ is $\Delta_{r}(a)$, thus, repeating such a process, we have
$[a]=s_{\phi}(a)+Vs_{\phi}(\Delta_{1}(a))+\cdots+V^{n-1}s_{\phi}(\Delta_{n-1}(a))+V^{n}\Delta_{n}(a).$
Furthermore, considering the ghost component $\varphi_{n}$, we have
(3.1) $a^{p^{n}}=\phi^{n}(a)+p\phi^{n-1}\Delta_{1}(a)+\cdots+p^{n}\Delta_{n}(a)=\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a).$
Moreover, we have
$p\Delta_{1}(a^{p^{n-1}})$ $=a^{p^{n}}-\phi(a^{p^{n-1}})$
=$\overset{(\star_{1}}{=}\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)-\sum^{n-1}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)$
$=p^{n}\Delta_{n}(a),$
where $(\star_{1})$ follows from (3.1). The last assertion follows from the proof of [kty]*Theorem 3.20. ∎
Theorem 3.7.
We use Notation 3.1. We regard $W_{n}(A)$ as $A$-module by $s_{\phi}$. Then we obtain an $A$-module isomorphism
$\Psi_{n}\colon W_{n}(A)\xrightarrow{\sim}A\oplus\phi_{*}A\oplus\cdots\oplus\phi^{n-1}_{*}A,$
where $\Psi_{n}$ is defined by
$\Psi_{n}((a_{0},\ldots,a_{n-1}))=(a_{0},\Delta_{1}(a_{0})+a_{1},\ldots,\Delta_{n-1}(a_{0})+\Delta_{n-2}(a_{1})+\cdots+\Delta_{1}(a_{n-2})+a_{n-1}).$
Proof.
For $n=1$, the assertion is clear, thus we assume $n\geq 2$. By Proposition 3.3, we obtain the exact sequence
[…]
of $A$-modules. Since $s_{\phi}\colon A\to W_{n}(A)$ is a splitting map of $R^{n-1}\colon W_{n}(A)\to A$, we obtain an $A$-module isomorphism
$W_{n}(A)\xrightarrow{\sim}\phi_{*}W_{n-1}(A)\oplus A\ ;\ \alpha\mapsto(\Delta_{W_{n}}(\alpha),R^{n-1}(\alpha)).$
Repeating to take decompositions, we obtain the isomorphism
$\Psi_{n}\colon W_{n}(A)\xrightarrow{\sim}A\oplus\phi_{*}A\oplus\cdots\oplus\phi^{n-1}_{*}A,$
and $\Psi_{n}$ satisfies
$\Psi_{n}((a_{0},\ldots,a_{n-1}))=(a_{0},\Delta_{1}(a_{0})+a_{1},\ldots,\Delta_{n-1}(a_{0})+\Delta_{n-2}(a_{1})+\cdots+\Delta_{1}(a_{n-2})+a_{n-1}),$
as desired. ∎
Proposition 3.8.
We use Notation 3.1. The $A$-module homomorphism $s_{\phi}\circ\phi\colon A\to\phi_{*}W_{n}(A)$ induces $\overline{s_{\phi}\circ\phi}\colon A/pA\to\phi_{*}W_{n}(A)/pW_{n}(A)$. Then $\overline{s_{\phi}\circ\phi}(\overline{a})=[a^{p}]$ for $a\in A$ and $\overline{a}$ is the image of $a$ by $A\to A/pA$.
Proof.
It is enough to show $\Psi_{n}(s_{\phi}\circ\phi(a))\equiv\Psi_{n}([a^{p}])\ \textrm{mod}\ p$ for $a\in A$. By Proposition 3.5 (1), we have $\Delta_{W_{n}}(s_{\phi}\circ\phi(a))=0$, thus $\Psi_{n}(s_{\phi}\circ\phi(a))=(\phi(a),0,\ldots,0)$. On the other hand, by Proposition 3.5 (3), we have
$\Delta_{W_{n}}([a^{p}])\equiv ps_{\phi}\circ\phi(a^{p-1})\Delta_{W_{n}}([a])\equiv 0\ \textrm{mod}\ pW_{n-1}(A).$
Thus, we have $\Psi_{n}([a^{p}])=(a^{p},0,\ldots,0)$. Since $\phi(a)\equiv a^{p}\ \textrm{mod}\ p$, we obtain $\Psi_{n}(s_{\phi}\circ\phi(a))\equiv\Psi_{n}([a^{p}])\ \textrm{mod}\ p$, as desired. ∎
4. Quasi-$F$-splitting in mixed characteristic
4.1. Definition of quasi-$F$-splitting
In this subsection, we introduce the notion of quasi-$F$-splitting for rings in mixed characteristic and study first properties of quasi-$F$-splitting.
Proposition 4.1.
Let $R$ be a $\mathbb{Z}_{(p)}$-algebra and $n$ a positive integer. We set $\overline{R}:=R/pR$ and $\overline{W}_{n}(R):=W_{n}(R)/pW_{n}(R)$. Then the map
$\Phi_{R,n}\colon\overline{R}\to\overline{W}_{n}(R)\ ;\ a\mapsto[\tilde{a}^{p}]$
is well-defined and a ring homomorphism, where $\tilde{a}\in R$ is a lift of $a\in\overline{R}$.
Proof.
Let $A:=\mathbb{Z}_{(p)}[X_{f}\mid f\in R]$ and $\phi\colon A\to A$ be the $\mathbb{Z}_{(p)}$-algebra satisfying $\phi(X_{f})=X^{p}_{f}$ for every $f\in R$. Let $\pi\colon A\to R$ be the $\mathbb{Z}_{(p)}$-algebra homomorphism defined by $\pi(X_{f})=f$. By Proposition 3.8, the map
$\Phi_{A,n}\colon\overline{A}:=A/pA\to\overline{W}_{n}(A):=W_{n}(A)/pW_{n}(A)\ ;\ a\mapsto[\tilde{a}^{p}]$
is well-defined and a ring homomorphism, where $\tilde{a}\in A$ is a lift of $a\in A$. We take $a,b\in A$ such that $\pi(a)=0$. Then we have $[(a+pb)^{p}]\equiv[a^{p}]\ \textrm{mod}\ p$ and $[a^{p}]$ is contained in the kernel of $W_{n}(\pi)$. Thus, $\Phi_{A,n}$ induces the ring homomorphism $\Phi_{R,n}$, as desired. ∎
Notation 4.2.
Let $R$ be a $\mathbb{Z}_{(p)}$-algebra. We set $\overline{R}:=R/pR$ and
$\overline{W}_{n}(R):=W_{n}(R)/pW_{n}(R)$
for each positive integer $n$. We defined an $\overline{R}$-module $Q_{R,n}$ by $Q_{R,n}:=\overline{W}_{n}(R)$ as group and the $\overline{R}$-module structure is endowed by the ring homomorphism $\Phi_{R,n}$ for each positive integer $n$.
Proposition 4.3.
We use Notation 4.2. We assume $R$ is $p$-torsion free. For positive integers $n,m$, we obtain the following exact sequence as $\overline{R}$-modules:
$0\to F^{m}_{*}Q_{R,n}\xrightarrow{V^{m}}Q_{R,n+m}\xrightarrow{R^{n}}Q_{R,m}\to 0$
In particular, if $R$ is Noether and $\overline{R}$ is $F$-finite, then $Q_{R,n}$ is a finite $\overline{R}$-module for every positive integer $n$.
Proof.
First, let us show that $V^{m}\colon F^{m}_{*}Q_{R,n}\to Q_{R,n+m}$ is an $\overline{R}$-module homomorphism. We take $\alpha\in W_{n}(R)$ and $b\in R$. Then we have $[b^{p}]V^{m}\alpha=V([b^{p^{m+1}}]\alpha)$ by Proposition 2.3, thus $V^{m}$ is an $\overline{R}$-module homomorphism. Next, let us show the exactness. It is enough to show that the sequence
$0\to\overline{W}_{n}(R)\xrightarrow{V^{m}}\overline{W}_{n+m}(R)\xrightarrow{R^{n}}\overline{W}_{m}(R)\to 0$
is exact as groups. It follows from the facts that $W_{s}(R)$ is $p$-torsion free for each positive integer $s$ and we have the exact sequence
$0\to W_{n}(R)\xrightarrow{V^{m}}W_{n+m}(R)\xrightarrow{R^{n}}W_{m}(R)\to 0$
as groups. The last assertion follows from the exact sequence and $Q_{R,1}=F_{*}\overline{R}$. ∎
Definition 4.4.
(cf. [Yobuko19]) We use Notation 4.2. We assume $R$ is Noether and $\overline{R}$ is $F$-finite.
  • Let $n$ be a positive integer. We say that $R$ is $n$-quasi-$F$-split if $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$ splits as $\overline{R}$-modules. Furthermore, we say that $R$ is quasi-$F$-split if $R$ is $n$-quasi-$F$-split for some positive integer $n$.
  • The quasi-$F$-splitting height $\mathrm{ht}(R)$ is defined by
[…]
    if $R$ is quasi-$F$-split, and $\mathrm{ht}(R):=\infty$ if $R$ is not quasi-$F$-split.
  • Let $n$ be a positive integer and $f\in R$ a non-zero divisor. We define an $\overline{R}$-module $Q^{f}_{R,n}$ by
    $Q^{f}_{R,n}:=\operatorname{Ker}(Q_{R,n}\to Q_{R/f,n}).$
    We note that
    $Q^{f}_{R,n}=\{(fa_{0},\ldots,fa_{n-1})\in\overline{W}_{n}(R)\mid a_{0},\ldots a_{n-1}\in R\}$
    as sets. We say that $(R,f)$ is purely $n$-quasi-$F$-split if the map
    $\Phi^{f}_{R,n}\colon\overline{R}\to Q^{f}_{R,n}\ ;\ a\mapsto[af]^{p}$
    splits as $\overline{R}$-modules.
  • Let $n,e$ be positive integers. We define $\overline{R}$-module $Q_{R,n,e}$ by the following pushout diagram as $\overline{R}$-modules:
[…]
    and the induced $\overline{R}$-module homomorphism is denoted by
    $\Phi_{R,n,e}\colon R\xrightarrow{\Phi_{R,n}}Q_{R,n}\to Q_{R,n,e}.$
  • We say that $R$ is quasi-$(F,F^{\infty})$-split if $R$ is quasi-$F$-split and $\Phi_{R,\mathrm{ht}(R),e}$ splits as $\overline{R}$-modules for every positive integer $e$.
Remark 4.5.
If $R$ is an $\mathbb{F}_{p}$-algebra, then the notion of quasi-$F$-splitting coincides with that in [Yobuko19].
Proposition 4.6.
We use Notation 4.2. We assume $R$ is Noetherian $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\mathrm{ht}(R)\leq\mathrm{ht}(\overline{R})$. Furthermore, if $R$ is normal and $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})=\operatorname{ht}^{e}(\overline{R})$ for every $e\geq 1$, where $\operatorname{ht}^{e}(\overline{R})$ is defined in [TWY]*Definition 3.5, then $R$ is quasi-$(F,F^{\infty})$-split.
Proof.
The first assertion follows from the following diagram:
[…]
for each positive integer $n$. We set $n:=\operatorname{ht}(R)$. Since we have the commutative diagram
[…]
we obtain the $\overline{R}$-module homomorphism $Q_{R,n,e}\to Q^{e}_{\overline{R},n}$, and in particular, we have
$\operatorname{Hom}_{W_{n}(\overline{R})}(Q^{e}_{\overline{R},n},W_{n}\omega_{\overline{R}}(-K_{\overline{R}}))$ $\to\operatorname{Hom}_{W_{n}(\overline{R})}(Q_{R,n,e},W_{n}\omega_{\overline{R}}(-K_{\overline{R}}))$
$\simeq\operatorname{Hom}_{\overline{R}}(Q_{R,n,e},\overline{R})\to\operatorname{Hom}_{\overline{R}}(\overline{R},\overline{R})\simeq\overline{R}.$
Since the composition is surjective by $\operatorname{ht}^{e}(\overline{R})=n$, the homomorphism $\overline{R}\to Q_{R,n,e}$ splits, as desired. ∎
Proposition 4.7.
We use Notation 4.2 and assume $(R,\mathfrak{m})$ is Noether local $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\operatorname{ht}(R)=\operatorname{ht}(\widehat{R})$, where $\widehat{R}$ is the $m$-adic completion of $R$. Furthermore, $R$ is quasi-$(F,F^{\infty})$-split if and only if so is $\widehat{R}$.
Proof.
First, we obtain $F^{e}_{*}Q_{R,n}\otimes_{R}\widehat{R}\simeq F^{e}_{*}Q_{\widehat{R},n}$ for positive integers $e,n$ by Proposition 4.3 and an inductive argument. Therefore, we obtain the first assertion. Furthermore, since taking pushout does not change the cokernels, we obtain the exact sequence
$0\to Q_{R,n}\to Q_{R,n,e}\to F^{e}_{*}B^{e-1}_{\overline{R}}\to 0,$
thus we obtain $Q_{R,n,e}\otimes_{R}\widehat{R}\simeq Q_{\widehat{R},e,n}$, and in particular, we have the last assertion. ∎
Proposition 4.8.
We use Notation 4.2 and assume $R$ is $p$-torsion free. If $R$ is Cohen-Macaulay, then so is $Q_{R,n}$ for every positive integer $n$.
Proof.
Let us show the assertion by induction on $n$. Since $R$ is Cohen-Macaulay and $p$-torsion free, the ring $\overline{R}\simeq W_{R,1}$ is Cohen-Macaulay. We take an integer $n\geq 2$. By Proposition 4.3, we obtain an exact sequence
$0\to F_{*}Q_{R,n-1}\to Q_{R,n}\to\overline{R}\to 0$
as $\overline{R}$-modules. By the induction hypothesis, the modules $F_{*}Q_{R,n-1}$ and $\overline{R}$ are Cohen-Macaulay, thus so is $Q_{R,n}$, as desired. ∎
Theorem 4.9.
We use Notation 4.2 and assume $(R,\mathfrak{m})$ is local. Let $f\in\mathfrak{m}$ be a regular element such that $R/f$ is $p$-torsion free. We assume $R$ is Gorenstein. Then $(R,f)$ is purely quasi-$F$-split if and only if $R/f$ is quasi-$F$-split.
Proof.
We obtain the following diagram in which each horizontal sequence is exact:
[…]
where $\overline{R/f}:=R/(f,p)$. Taking local cohomology we obtain the following diagram in which each horizontal sequence is exact:
[…]
where $d:=\dim\overline{R}$ and the injectivity of ()($\star$)( ⋆ ) follows from Proposition 4.8. Since the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ is contained in the image of the map $H^{d-1}_{\mathfrak{m}}(\overline{R/f})\to H^{d}_{\mathfrak{m}}(\overline{R})$, we obtain that $\Phi_{R/f,n}$ is injective if and only if so is $\Phi^{f}_{R,n}$, as desired. ∎
4.2. Quasi-$F$-splitting for perfectoids
In this subsection, we study the purity of $\Phi_{R,n}$ for a perfectoid $R$. Theorem 4.10 is related to the computation of perfectoid pure threshold (cf. Theorem 5.8).
Theorem 4.10.
Let $R$ be a $p$-torsion free perfectoid and $n$ a positive integer. We assume $R$ has a compatible system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. Then there exist an $\overline{R}$-module homomorphism $\overline{\psi}_{n}\colon Q_{R,n}\to R/(p^{i_{n}})$ and a unit $w\in\overline{R}$ such that $\overline{\psi}_{n}(1)=1$ and $\overline{\psi}_{n}(V^{n-1}[a])=wp^{i_{n-1}}F^{-n}(a)$ for every $a\in R$, where
$i_{m}:=\frac{1}{p}+\cdots+\frac{1}{p^{m}}$
for every integer $m\geq 1$ and $i_{0}:=0$.
Proof.
Let $R^{\flat}$ be a tilt of $R$ and $T\in R^{\flat}$ the element corresponding to a system $\{p^{1/p^{e}}\}$. Let $A:=W(R^{\flat})$ and $d:=[T]-p$, then $(A,(d))$ is a perfect prism with $A/(d)\simeq R$ by [BS]*Corollary 2.31. The lift of Frobenius on $A$ is denoted by $\phi$. Let us show the following claim:
Claim 4.11.
There exists an $A$-module homomorphism $\psi_{n}\colon Q_{A,n}\to R^{\flat}$ such that
  1. (a)
    $\psi_{n}(1)$ and $\psi_{n}(\Delta_{W_{n+1}}(d))$ are units,
  2. (b)
    $\psi_{n}(W_{n}((d))\cdot\overline{W}_{n}(A))\subseteq(T^{i_{n}})$, and
  3. (c)
    there exists a unit $w\in R^{\flat}$, $\psi_{n}(V^{n-1}[a])=wT^{i_{n-1}}F^{-n}(a)$ for every $a\in A$.
Proof.
We prove the assertions by induction on $n$. For $n=1$, we define $\psi_{1}$ by
$\psi_{1}\colon Q_{A,1}\simeq F_{*}R^{\flat}\xrightarrow{F^{-1}}R^{\flat},$
then $\psi_{1}(1)=1$ and
$\psi_{1}(\Delta_{W_{2}}(d))=\psi_{1}(\delta_{1}(d))=F^{-1}(\delta_{1}(d)).$
Since $\delta_{1}(d)$ is a unit, the map $\psi_{1}$ satisfies the condition (a). On the other hand, we have
$\psi_{1}(W_{1}((d))\cdot\overline{W}_{1}(A))=F^{-1}(TR^{\flat})=T^{1/p}R^{\flat},$
thus $\psi_{1}$ also satisfies the condition (b). Furthermore, we have $\psi(a)=F^{-1}(a)$, thus $\psi$ satisfies the condition (c). Next, we assume $n\geq 2$. By the induction hypothesis, there exists $\psi_{n-1}\colon Q_{A,n-1}\to R^{\flat}$ such that $\psi_{n-1}$ satisfies the conditions (a), (b), (c). We replace $\psi_{n-1}$ by $(-\psi_{n-1}(\Delta_{W_{n}}(d))^{-1})\cdot\psi_{n-1}$, we may assume that $\psi_{n-1}(\Delta_{W_{n}}(d))=-1$. We set
$\psi^{\prime}:=T^{1/p}\cdot F^{-1}\circ\psi_{n-1}\colon F_{*}Q_{A,n-1}\to R^{\flat}.$
By the proof of Theorem 3.7, we obtain an isomorphism
$Q_{A,n}\xrightarrow{\sim}F_{*}Q_{A,n-1}\oplus F_{*}R^{\flat}\ ;\ \alpha\mapsto(\Delta(\alpha),R^{n-1}(\alpha)),$
and we define an $A$-module homomorphism $\psi_{n}$ by
$\psi_{n}\colon\phi_{*}\overline{W}_{n}(A)\to\phi^{2}_{*}W_{n-1}(A)\oplus F_{*}R^{\flat}\xrightarrow{(\psi^{\prime},F^{-1})}R^{\flat},$
that is, for $\alpha=(a_{0},\ldots,a_{n-1})\in\phi_{*}\overline{W}_{n}(A)$,
$\psi_{n}(\alpha)=\psi^{\prime}(\Delta_{W_{n}}(\alpha))+F^{-1}(a_{0}).$
We take $a_{0},\ldots,a_{n-1}\in A$, we have
$\psi_{n}((a_{0}d,\ldots,a_{n-1}d))$
$=$ $\psi^{\prime}(\Delta_{W_{n}}([a_{0}d])+(a_{1}d,\ldots,a_{n-1}d))+F^{-1}(a_{0}d)$
$=$ $F^{-1}(a_{0})(\psi^{\prime}(\Delta_{W_{n}}(d))+T^{1/p})+\psi^{\prime}([d^{p}]\Delta_{W_{n}}([a_{0}])+(a_{1}d,\ldots,a_{n-1}d))$
=$\overset{(\star_{1})}{=}$ $T^{1/p}\psi^{\prime}(\Delta_{W_{n}}([a_{0}]))+\psi^{\prime}((a_{1}d,\ldots,a_{n-1}d))\overset{(\star_{2})}{\in}(T^{i_{n}}),$
where $(\star_{1})$ follows from
$\psi^{\prime}(\Delta_{W_{n}}(d))+T^{1/p})=T^{1/p}(F^{-1}\circ\psi_{n-1}(\Delta_{W_{n}}(d))+1)=0$
and $(\star_{2})$ follows from the facts
$T^{1/p}\psi^{\prime}(\Delta_{W_{n}}([a_{0}]))\in(T^{2/p})\subseteq(T^{i_{n}})$
and
$\psi^{\prime}((a_{1}d,\ldots,a_{n-1}d))$ $=T^{1/p}F^{-1}\circ\psi_{n-1}((a_{1}d,\ldots,a_{n-1}d))$
$\subseteq T^{1/p}\cdot F^{-1}((T^{i_{n-1}}))=(T^{i_{n}}).$
Thus, the homomorphism $\psi_{n}$ satisfies the condition (b). Furthermore, we have $\psi_{n}(1)=F^{-1}(1)=1$ and
$\psi_{n}(\Delta_{W_{n+1}}(d))$ $=\psi^{\prime}(\Delta_{W_{n}}\circ\Delta_{W_{n+1}}(d))+F^{-1}(\delta_{1}(d))$
$\equiv F^{-1}(\delta_{1}(d))\ \textrm{mod}\ T^{1/p}.$
Since $R^{\flat}$ is $T^{1/p}$-complete and $F^{-1}(\delta_{1}(d))$ is a unit, the element $\psi_{n}(\Delta_{W_{n+1}}(d))$ is a unit. Furthermore, we have
$\psi(V^{n-1}[a])=\psi^{\prime}(V^{n-2}[a])=T^{1/p}\cdot F^{-1}(wT^{i_{n-2}}F^{-(n-1)}(a))=F^{-1}(w)T^{i_{n-1}}F^{-n}(a),$
thus, the map $\psi_{n}$ satisfies the conditions (a), (b), (c). $\blacksquare$
We take $\psi_{n}$ as in Claim 4.11. Replacing $\psi_{n}$ by $(\psi_{n}(1))^{-1}\cdot\psi_{n}$, we may assume $\psi_{n}(1)=1$. By the condition (b), $\psi_{n}$ induces the map
$\overline{\psi}_{n}\colon Q_{R,n}\to R^{\flat}/(T^{i_{n}})\simeq R/(p^{i_{n}})$
with $\overline{\psi}_{n}(1)=1$. Furthermore, by the condition (c), there exists a unit $w\in\overline{R}$ such that $\overline{\psi}(V^{n-1}[a])=wT^{i_{n}}F^{-n}(a)$ for every $a\in A$. ∎
4.3. Fedder type criterion for quasi-$F$-splitting
In this subsection, we introduce Fedder type criterion (Theorem 4.12) for quasi-$F$-splitting and quasi-$(F,F^{\infty})$-splitting. The proof of Theorem 4.12 is almost identical to a proof of [TWY]*Theorem A, however, for the convenience of the reader, we provide a proof in Appendix C. Furthermore, we introduce examples of computations of quasi-$F$-splitting height and quasi-$(F,F^{\infty})$-splitting by using Theorem 4.12.
Theorem 4.12.
(cf. [kty]*Theorem A) Let $(A,\mathfrak{m})$ be a regular local ring with lift of Frobenius $\phi\colon A\to A$ such that $\phi$ is finite. We assume $p\in\mathfrak{m}\backslash\mathfrak{m}^{2}$ and set $\overline{A}:=A/pA$, then $\overline{A}$ is regular. We fix a generator $u\in\operatorname{Hom}_{\overline{A}}(F_{*}\overline{A},\overline{A})$ as $F_{*}\overline{A}$-module. Let $f_{1},\ldots,f_{r}$ be a regular sequence in $A$ and we set $I:=(f_{1},\ldots,f_{r})\overline{A}$, $f:=f_{1}\cdots f_{r}$. We assume $R:=A/(f_{1},\ldots,f_{r})$ is $p$-torsion free.
  1. (1)
    We define a sequence $\{I_{n}\}$ of ideals of $\overline{A}$ by $I_{1}:=f^{p-1}\overline{A}+I^{[p]}$ and
    $I_{n}:=u(F_{*}(\Delta_{1}(f^{p-1})I_{n-1}))+I_{1}$
    for $n\geq 2$, inductively. Then we have
    $\operatorname{ht}(R)=\inf\{n\geq 1\mid I_{n}\nsubseteq\mathfrak{m}^{[p]}\},$
    where if $I_{n}\subseteq\mathfrak{m}^{[p]}$ for every positive integer $n$, then the right hand side is defined by $\infty$.
  2. (2)
    We set $I^{\prime}:=\bigcap_{e\geq 1}u^{e}(F^{e}_{*}(f^{p^{e}-1}\overline{A}))$. We define a sequence $\{I^{\prime}_{n}\}$ of ideals of $\overline{A}$ by $I^{\prime}_{1}:=f^{p-1}I^{\prime}+I^{[p]}$ and
    $I^{\prime}_{n}:=u(F_{*}(\Delta_{1}(f^{p-1})I^{\prime}_{n-1}))+f^{p-1}\overline{A}+I^{[p]}$
    for $n\geq 2$, inductively. Then $R$ is quasi-$(F,F^{\infty})$-split if and only if $R$ is quasi-$F$-split and $I^{\prime}_{\operatorname{ht}(R)}\nsubseteq\mathfrak{m}^{[p]}$.
Example 4.13.
(cf. Example 5.9) Let $A:=\mathbb{Z}_{(p)}[[X_{1},\ldots,X_{N}]]$ and the $\mathbb{Z}_{(p)}$-algebra homomorphism $\phi$ is given by $\phi(X_{i})=X_{i}^{p}$. If $N=3$, we set $x:=X_{1},y:=X_{2},z:=X_{3}$, and if $N=4$ we set $x:=X_{1},y:=X_{2},z:=X_{3},w:=X_{4}$. Then $F_{*}\overline{A}$ has a basis $\{F_{*}(X_{1}^{i_{1}}\cdots X_{N}^{i_{N}})\mid p-1\geq i_{1},\ldots,i_{N}\geq 0\}$. The corresponding dual basis is denoted by $u_{1},\ldots,u_{d}$ and the corresponding element to $X_{1}^{p-1}\cdots X_{N}^{p-1}$ is denoted by $u$, then $u$ is a generator of $\operatorname{Hom}_{\overline{A}}(F_{*}\overline{A},\overline{A})$ as $F_{*}\overline{A}$-module. Let $f\in A$ be a nonzero divisor such that $R:=A/f$ is $p$-torsion free. We use same notations $I_{n},I^{\prime}_{n},I^{\prime}$ as in Theorem 4.12 for every positive integer $n$. We set $\overline{R}:=R/pR$ and define $\theta\colon F_{*}\overline{A}\to\overline{A}$ by $\theta(F_{*}a)=u(F_{*}(a\Delta_{1}(f)))$.
  1. (1)
    Let $p=2,N=3,f=x^{3}+y^{3}+z^{3}$. Then $I_{1}\subseteq\mathfrak{m}^{[p]}$ and $u(F_{*}(f\Delta_{1}(f))\equiv xyz\ \textrm{mod}\ \mathfrak{m}^{[p]}$, thus we have $I_{2}\nsubseteq\mathfrak{m}^{[p]}$. Furthermore, we have $u(F_{*}(f\Delta_{1}(f)\mathfrak{m}))\subseteq\mathfrak{m}^{[p]}$, thus we have $I_{2}^{\prime}\subseteq\mathfrak{m}^{[p]}$, in conclusion, $R$ is not quasi-$(F,F^{\infty})$-split (cf. Proposition 6.8).
  2. (2)
    Let $N=3,f=z^{2}+x^{3}+y^{5}$.
    • We consider the case of $p=2$, then we have
      (4.1) ${zf}$${xy^{2}z}$${y^{3}z}$${xyz\notin\mathfrak{m}^{[p]},}$$\scriptstyle{\theta}$$\scriptstyle{\theta}$$\scriptstyle{\theta}$
      thus we have $\operatorname{ht}(R)\leq 4$. Furthermore, since we have $f\Delta_{1}(f)^{3}\in\mathfrak{m}^{[p^{3}]}$, thus we have $I_{3}\subseteq\mathfrak{m}^{[p]}$, therefore $\operatorname{ht}(R)=4$. Moreover, since $I^{\prime}=(x,y^{2},z)$ and the first element $zf$ of (4.1) is contained in $I_{1}^{\prime}$, the ring $R$ is quasi-$(F,F^{\infty})$-split.
    • We consider the case of $p=3$, then we have
      ${x^{2}yf^{2}}$${-x^{2}y^{3}f}$${x^{2}y^{2}f\equiv x^{2}y^{2}z^{2}\notin\mathfrak{m}^{[p]},}$$\scriptstyle{\theta}$$\scriptstyle{\theta}$
      $f^{2}\Delta_{1}(f^{2})\in\mathfrak{m}^{[p^{2}]}$, and $I^{\prime}=(x,y^{3},z)$, thus we have $\operatorname{ht}(R)=3$ and $R$ is quasi-$(F,F^{\infty})$-split.
    • We consider the case of $p=5$, then we have
      ${y^{4}f^{4}}$${xy^{4}f^{3}\equiv-2x^{4}y^{4}z^{4}\notin\mathfrak{m}^{[p]},}$$\scriptstyle{\theta}$
      $f^{4}\in\mathfrak{m}^{[p]}$, and $I^{\prime}=(x,y,z)$, thus we have $\operatorname{ht}(R)=2$ and $R$ is quasi-$(F,F^{\infty})$-split.
  3. (3)
    Let $p=2,N=4,f=w^{2}+xyz(x+y+z)$. Then $I_{1}\subseteq\mathfrak{m}^{[p]}$ and
    $u(F_{*}(zwf\Delta_{1}(f)))=xyzw\notin\mathfrak{m}^{[p]},$
    thus we have $\operatorname{ht}(R)=2$. We note that $\operatorname{ht}(\overline{R})=\infty$ by [kty]*Example 7.11. Furthermore, the ring $R$ is quasi-$(F,F^{\infty})$-split by $zw\in I^{\prime}$.
  4. (4)
    Let $p=2,N=4,f=w^{2}+xyz(x+y+z)+p(xy+xz+yz)w$. Then $\operatorname{ht}(R)=3$. Indeed,
    $I_{2}$ $=$ $(f,zw^{2}+x^{2}yz^{2}+xy^{2}z^{2}+xz^{2}w+yz^{2}w,xw^{2}+x^{2}y^{2}z+x^{2}yz^{2}+x^{2}yw+x^{2}zw,$
    $xyw^{2}+xzw^{2}+yzw^{2}+w^{3}+x^{2}y^{2}z^{2})\subseteq\mathfrak{m}^{[p]},$
    and
    $\theta(F_{*}(xw^{2}+x^{2}y^{2}z+x^{2}yz^{2}+x^{2}yw+x^{2}zw))=xzw+xyw\notin\mathfrak{m}^{[p]},$
    thus we have $\operatorname{ht}(R)=3$. On the other hand, we have
    ${xwf\Delta_{1}(f)}$${xw^{2}+x^{2}y^{2}z+x^{2}yz^{2}+x^{2}yw+x^{2}zw}$${xzw+xyw}$$\scriptstyle{\theta}$$\scriptstyle{\theta}$
    and $xw\in I^{\prime}$, thus $R$ is quasi-$(F,F^{\infty})$-split.
  5. (5)
    Let $p=2,N=3,f=z^{2}+x^{2}y+xy^{n}$ for $n\geq 2$. First, we note that $z\in I^{\prime}$. By the argument in [kty]*Section 7.1, we have $\theta^{m}(F^{m}_{*}(zy^{2^{m}-n}f))=xyz\notin\mathfrak{m}^{[p]}$ and $\operatorname{ht}(R)=m+1$, where $m:=\lceil\log_{2}n\rceil$, thus $R$ is quasi-$(F,F^{\infty})$-split.
Example 4.14.
Let $A:=\mathbb{Z}_{(p)}[[x,y,z,x^{\prime},y^{\prime},z^{\prime}]]$, $f:=x^{3}+y^{3}+z^{3}$, $f^{\prime}:=x^{\prime 3}+y^{\prime 3}+z^{\prime 3}$, $R:=A/(f,f^{\prime})$, and $p\equiv 2\ \textrm{mod}\ 3$. Then $R$ is not quasi-$F$-split. Indeed, we have
$\Delta_{1}((ff^{\prime})^{p-1})=f^{p(p-1)}\Delta_{1}(f^{\prime p-1})+f^{\prime p(p-1)}\Delta_{1}(f^{p-1})\in\mathfrak{m}^{[p^{2}]},$
thus we have $I_{n}\subseteq\mathfrak{m}^{[p]}$ for every integer $n\geq 1$.
5. Quasi-$F$-splitting versus perfectoid purity
In this section, we study the relationship between quasi-$F$-splitting and perfectoid purity. For this purpose, we introduce a program (Definition 5.4) to obtain a sequence of elements of local cohomologies. Whether or not this program stops is closely related to the singularity of a given ring.
Notation 5.1.
Let $(R,\mathfrak{m})$ be a Noetherian local $\mathbb{Z}_{(p)}$-algebra such that $R$ is $p$-torsion free, $p\in\mathfrak{m}$, and $\overline{R}:=R/pR$ is $F$-finite. We set $d:=\dim{\overline{R}}$.
Proposition 5.2.
We use Notation 5.1. Let $n$ be a positive integer and $i_{n}=1/p+\cdots+1/p^{n}$. We assume $R$ is Gorenstein. Let $\rho\colon R\to R_{\infty}$ be a ring homomorphism to a $p$-torsion free perfectoid $R_{\infty}$. If $\operatorname{ht}(R)\geq n+1$, then
$H^{d+1}_{\mathfrak{m}}(R)\to H^{d+1}_{\mathfrak{m}}(R_{\infty})\xrightarrow{\cdot p^{1-i_{n}}}H^{d+1}_{\mathfrak{m}}(R_{\infty})$
is not injective. In particular, we have $\operatorname{ppt}(R;{\rm div}(p))\leq 1-i_{n}$.
Proof.
We consider the following commutative diagram
${H^{d}_{\mathfrak{m}}(\overline{R})}$${H^{d}_{\mathfrak{m}}(Q_{R,n})}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})}$${H^{d}_{\mathfrak{m}}(Q_{R_{\infty},n})}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{i_{n}})),}$$\scriptstyle{H^{d}_{\mathfrak{m}}(\rho)}$$\scriptstyle{H^{d}_{\mathfrak{m}}(\overline{\psi}_{n})}$
where $\overline{\psi}_{n}$ is as in Theorem 4.10. In particular, taking a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$, since $\eta$ is zero in $H^{d}_{\mathfrak{m}}(Q_{R,n})$ by $\operatorname{ht}(R)\geq n+1$, we have
$H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\rightarrow H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{i_{n}}))\ ;\ \eta\mapsto\rho(\eta)\mapsto 0.$
By the commutative diagram in which each horizontal sequence is exact;
[…]
we obtain the commutative diagram
${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})}$${H^{d+1}_{\mathfrak{m}}(R_{\infty})}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{i_{n}}))}$${H^{d+1}_{\mathfrak{m}}(R_{\infty}).}$$\scriptstyle{\cdot p^{1-i_{n}}}$
Therefore, by the homomorphism
(5.1) $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(R)\xrightarrow{H^{d+1}_{\mathfrak{m}}(\rho)}H^{d+1}_{\mathfrak{m}}(R_{\infty})\xrightarrow{\cdot p^{1-i_{n}}}H^{d+1}_{\mathfrak{m}}(R_{\infty})$
the element $\eta$ maps to 00. Since $R$ is $p$-torsion free and Gorenstein, the first homomorphism in (5.1) is injective, thus we obtain that the map
$H^{d+1}_{\mathfrak{m}}(R)\to H^{d+1}_{\mathfrak{m}}(R_{\infty})\xrightarrow{\cdot p^{1-i_{n}}}H^{d+1}_{\mathfrak{m}}(R_{\infty})$
is not injective, as desired. The last assertion follows from the existence of $p$-torsion free test perfectoids (Theorem A.1) and Proposition 2.9. ∎
Corollary 5.3.
We use Notation 5.1. We assume $R$ is Gorenstein. If $R$ is not quasi-$F$-split and $p=2$, then $R$ is not perfectoid BCM-regular.
Proof.
If $R$ is perfectoid BCM-regular, then there exists $\varepsilon\in\mathbb{Z}[\frac{1}{p}]_{>0}$ such that $(R,\varepsilon{\rm div}(p))$ is perfectoid BCM-regular by [bcm-reg]*Proposition 6.10. Thus, there exists a perfectoid big Cohen-Macaulay algebra $B$ over $R$ such that $R\to B\xrightarrow{\cdot p^{\varepsilon}}B$ is pure. Since $B$ is $p$-torsion free, it contradicts the assumption by Proposition 5.2. ∎
Definition 5.4.
We use Notation 5.1. We take $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and consider the following program. We set $\eta_{0}:=\eta$ and assume $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ is defined for $i\geq 0$.
  1. Case(i):
    We consider the case of $F(\eta_{i})\neq 0$, where $F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(F_{*}\overline{R})$.
    1. Case(i-i):
      If there exists $e^{\prime}\geq 2$ such that $F^{e^{\prime}}(\eta_{i})=0$, then we take the integer $e\geq 1$ such that $F^{e}(\eta_{i})\neq 0$ and $F^{e+1}(\eta_{i})=0$ and we set $\eta_{i+1}:=F^{e}(\eta_{i})$.
    2. Case(i-ii):
      If $F^{e}(\eta_{i})\neq 0$ for every integer $e\geq 2$, then we stop the program.
  2. Case(ii):
    We consider the case of $F(\eta_{i})=0$.
    1. Case(ii-i):
      If there exists a positive integer $n^{\prime}\geq 2$ such that $\Phi_{R,n^{\prime}}(\eta)\neq 0$, where $\Phi_{R,n^{\prime}}\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(Q_{R,n^{\prime}})$, then we take an integer $n\geq 2$ such that $\Phi_{R,n-1}(\eta)=0$ and $\Phi_{R,n}(\eta)\neq 0$. By the exact sequence in Proposition 4.3, there exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})$ such that $V^{n-1}(\eta^{\prime})=\Phi_{R,n}(\eta_{i})$. Then we chose such $\eta^{\prime}$ and we set $\eta_{i+1}:=\eta^{\prime}$.
    2. Case(ii-ii):
      If $\Phi_{R,n}(\eta_{i})=0$ for every $n\geq 2$, then we stop the program.
Remark 5.5.
Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$.
  1. (1)
    In the program in Definition 5.4, if $\eta_{i}$ is in the case (i-i), then $\eta_{i+1}$ is in the case (ii).
  2. (2)
    The program in Definition 5.4 has the following three outputs.
    • We can define $\eta_{0},\eta_{1},\ldots,\eta_{i}$ for some $i\geq 0$ and $\eta_{i}$ is in the case (i-ii).
    • We can define $\eta_{0},\eta_{1},\ldots,\eta_{i}$ for some $i\geq 0$ and $\eta_{i}$ is in the case (ii-ii).
    • We can define $\eta_{i}$ for all $i\geq 0$.
  3. (3)
    If $R$ is quasi-$F$-split, the case (ii-ii) does not occur since $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(Q_{R,n})$ is injective for some $n\geq 1$.
Lemma 5.6.
We use Notation 5.1. Let $S$ be an $R$-algebra and $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$. We assume $S$ is $p$-torsion free and $H^{d-1}_{\mathfrak{m}}(\overline{S})=0$, where $\overline{S}:=S/pS$. We run a program in Definition 5.4 for $\eta_{0}:=\eta$ and get $\eta_{i}$ in the $i$-th step. If $\eta$ is contained in $\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{S}))$, then so is $\eta_{i}$.
Proof.
It is enough to show $\eta_{1}\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{S}))$. If $\eta$ is in the case (i-i), then $\eta_{1}=F^{e}(\eta)$ for some $e\geq 1$ and the assertion follows from the commutative diagram
[…]
If $\eta$ is in the case (ii-i), then $\Phi_{R,n}(\eta)=V^{n-1}(\eta_{1})$ for some $n\geq 1$. By the commutative diagram
[…]
it is enough to show $V^{n-1}\colon H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{S})\to H^{d}_{\mathfrak{m}}(Q_{S,n})$ is injective. By the exact sequence in Proposition 4.3 and $H^{d-1}_{\mathfrak{m}}(\overline{S})=0$, we obtain $H^{d-1}_{\mathfrak{m}}(Q_{S,n-1})=0$. Therefore, we have $V^{n-1}\colon H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{S})\to H^{d}_{\mathfrak{m}}(Q_{S,n})$ is injective by Proposition 4.3. ∎
Proposition 5.7.
We use Notation 5.1. We assume $R$ is complete intersection. We take a surjection $S\to R$ and a perfectoid $R_{\infty}:=R^{S_{\infty}}_{perfd}$ as in [p-pure]*Lemma 4.23. We run a program in Definition 5.4 for an element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and assume that we obtain $\eta:=\eta_{0},\ldots,\eta_{i}$ for some $i\geq 0$ and $\eta_{i}$ is in the case of (i-ii). Then $\eta\notin\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))$.
Proof.
We set $\rho\colon R\to R_{\infty}$. We note that $H^{d-1}_{\mathfrak{m}}(\overline{R}_{\infty})=0$ by [p-pure]*Theorem 4.24. Suppose $\eta\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{\rho}))$, where $\overline{\rho}\colon\overline{R}\to\overline{R}_{\infty}$. Then $\eta_{i}\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{\rho}))$ by Lemma 5.6. Since we have $(\overline{R}_{\infty})_{perfd}\simeq\overline{R}_{perfd}$, there exists a positive integer $e^{\prime}$ such that $F^{e^{\prime}}(\eta_{i})=0$, it contradicts the fact that $\eta_{i}$ is in the case (i-ii). ∎
Theorem 5.8.
(cf. Theorem A) We use Notation 5.1. We assume $R$ is complete intersection. If $R$ is quasi-$(F,F^{\infty})$-split and $\operatorname{ht}(R)=n$, then $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, where $i_{n-1}:=1/p+\cdots+1/p^{n-1}$.
Proof.
By [p-pure]*Lemma 4.8 and Proposition 4.7, we may assume that $R$ is $\mathfrak{m}$-adic complete. We take a surjection $S\to R$ and a perfectoid $R_{\infty}:=R^{S_{\infty}}_{perfd}$ be as in [p-pure]*Lemma 4.23. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a non-zero element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ and we run a program in Definition 5.4. If $\eta$ is in the case (i), then $\overline{R}$ is $F$-pure, and in particular, $\eta$ is in the case (i-ii). Thus, we obtain $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}((\overline{R}_{\infty})_{perf})$ is injective by the proof of Proposition 5.7. Since $(\overline{R}_{\infty})_{perf}\simeq R_{\infty}/(p^{1/p^{\infty}})$, $R$ is Gorenstein, and $H^{d}_{\mathfrak{m}}(R_{\infty})=0$, we have $(R,\varepsilon{\rm div}(p))$ is perfectoid pure for each $\varepsilon\in\mathbb{Z}[1/p]$ with $1>\varepsilon\geq 0$. In particular, we have $\operatorname{ppt}(R;{\rm div}(p))=1$. Next, we assume $\eta$ is in the case (ii), that is, $F(\eta)=0$, and in particular, we have $\operatorname{ht}(R)=n\geq 2$. Since $\eta$ is an element of the socle, we have $\Phi_{R,n}(\eta)\neq 0$ and $\Phi_{R,n-1}(\eta)=0$. Therefore, we obtain $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $V^{n-1}(F^{n}_{*}\eta_{1})=\Phi_{R,n}(\eta)$. If there exists a positive integer $e$ such that $F^{e}(\eta_{1})=0$, then $\eta$ is in the kernel of the map
$\Phi_{R,n,e}\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(Q_{R,n})\to H^{d}_{\mathfrak{m}}(Q_{R,n,e}),$
thus it contradicts the fact that $R$ is quasi-$(F,F^{\infty})$-split. Therefore, we obtain that $F^{e}(\eta_{1})\neq 0$ for every positive integer $e$, that is, $\eta_{1}$ is in the case of (i-ii), then by Lemma 5.6 and Proposition 5.7, we have $\eta$ is not contained in $\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))$, therefore by the above argument, $R$ is perfectoid pure. Next, let us compute $\operatorname{ppt}(R;{\rm div}(p))$. We take $\overline{\psi}_{n}\colon Q_{R_{\infty},n}\to\overline{R}_{\infty}$ and a unit $w\in\overline{R}_{\infty}$ as in Theorem 4.10. We consider the following commutative diagram;
[…]
where $\alpha$ follows from the natural surjection and the commutativity of right bottom diagram follows from Theorem 4.10. We consider the images of $\eta^{\prime}$ and $F^{-n}(\eta^{\prime}_{1})$ in $H^{d+1}_{\mathfrak{m}}(R_{\infty})$, then we obtain
(5.2) $p^{(1-i_{n})}(\eta^{\prime}/p)=wp^{i_{n-1}}F^{-n}(\eta^{\prime}_{1})/p^{i_{n}}=wF^{-n}(\eta^{\prime}_{1})/(p^{1/p^{n}}).$
Since $\eta^{\prime}_{1}$ is not contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{\infty}}))$ by $F^{e}(\eta)\neq 0$ for each $e\geq 1$, the element $F^{-n}(\eta^{\prime}_{1})$ is not contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{n}}))\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{\infty}})$, that is, we obtain $p^{\varepsilon}(F^{-n}(\eta^{\prime}_{1})/p^{1/p^{n}})\neq 0$ for every $1/p^{n}>\varepsilon\geq 0$. Therefore, by (5.2), we have $p^{\varepsilon}(\eta^{\prime}/p)=0$ if and only if $\varepsilon\geq 1-i_{n-1}$ for $\varepsilon\in\mathbb{Z}[\frac{1}{p}]$. In conclusion, we have $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$ by Proposition 2.9, as desired. ∎
Example 5.9.
By Theorem 5.8, we can compute perfectoid pure thresholds for examples in Example 4.13 other than (1). We will compute perfectoid pure threshold of Example 4.13 (1) in Proposition 6.10.
  1. (1)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z]]/(z^{2}+y^{3}+z^{5})$, then
    $\operatorname{ppt}(R;{\rm div}(p))=\begin{cases}1/8&\textup{if $p=2$},\\ 5/9&\textup{if $p=3$},\\ 4/5&\textup{if $p=5$},\\ 1&\textup{otherwise}.\end{cases}$
    We note that if $p\geq 7$, then $\overline{R}$ is $F$-pure.
  2. (2)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z,w]]/(w^{2}+xyz(x+y+z))$ and $p=2$, then $\operatorname{ppt}(R;{\rm div}(p))=1/2$.
  3. (3)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z,w]]/(w^{2}+xyz(x+y+z)+p(xy+xz+yz)w)$ and $p=2$, then $\operatorname{ppt}(R;{\rm div}(p))=1/4$.
  4. (4)
    Let $R:=\mathbb{Z}_{(p)}[[x,y,z]]/(z^{2}+x^{2}y+xy^{n})$ for $n\geq 2$ and $p=2$, then $\operatorname{ppt}(R;{\rm div}(p))=1/2^{\lceil\log_{2}n\rceil}$.
Theorem 5.10.
Let $(A,\mathfrak{m})$ be a local Gorenstein $\mathbb{Z}_{(p)}$-algebra such that $A$ is $p$-torsion free and $p\in\mathfrak{m}$. Let $f\in\mathfrak{m}$ be an element of $A$ such that $f,p$ is a regular sequence of $A$. Let $A_{\infty}$ be a perfectoid over $A$ such that $(A_{\infty}/(p))_{perfd}\simeq(A/(p))_{perf}$, $A_{\infty}$ is a cohomologically Cohen-Macaulay, $A\to A_{\infty}$ is pure, $A_{\infty}$ has a system of $p$-power roots system of $f$, and $f,p$ are regular elements of $A_{\infty}$. Let $R:=A/(f)$ and $R_{\infty}:=(A_{\infty}/(f))_{perfd}$. If $R$ is quasi-$F$-split, then $R\to R_{\infty}$ is pure.
Proof.
We set $\overline{A}:=A/(p)$, $\overline{R}:=R/(p)$, $\overline{R}_{\infty}:=R_{\infty}/(p)$, $\overline{A}_{\infty}:=A_{\infty}/(p)$, and $d:=\dim\overline{R}$. We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Suppose $R\to R_{\infty}$ is not perfectoid pure, then $\eta\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))$. We run a program in Definition 5.4 for $\eta$ and get $\eta_{0}:=\eta,\eta_{1},\eta_{2},\ldots$ by Lemmas 5.6 and 5.7. We fix a system of $p$-power roots $\{f^{1/p^{e}}\}$ of $f$ in $A_{\infty}$. By the exact sequence
$0\to\overline{A}\xrightarrow{\cdot f}\overline{A}\to\overline{R}\to 0,$
we obtain the homomorphism $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(\overline{A})$. We set
$\tau_{i}:=(u\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(\overline{A})\to H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty}))(\eta_{i})$
and $\tau_{i}^{(e)}:=\tau_{i}f^{1-1/p^{e}}$ for all integers $e,i\geq 0$. Let us show the following claim.
Claim 5.11.
We have $\tau_{i}^{(e)}\neq 0$ for all integers $e,i\geq 0$.
Proof.
Let us show the assertion by induction on $e$. Since $A$ is Cohen-Macaulay, the homomorphism $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(\overline{A})$ is injective. Since $A\to A_{\infty}$ is pure, the homomorphism $H^{d+1}_{\mathfrak{m}}(\overline{A})\to H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty})$ is injective, and in particular, we have $\tau_{i}^{(0)}=\tau_{i}\neq 0$ for every integer $i\geq 0$. Thus, we assume $e\geq 1$ and take $i\geq 0$. First, we assume that $\eta_{i}$ is in the case (i-i) in Definition 5.4. Then we have $F^{n}(\eta_{i})=\eta_{i+1}$ for some $n\geq 1$. By the commutative diagram
[…]
we have
$\tau_{i+1}^{(e-1)}$ $=u(\eta_{i+1})f^{1-1/p^{e-1}}=u\circ F^{n}(F^{n}_{*}\eta_{i})f^{1-1/p^{e-1}}$
$=F^{n}(F^{n}_{*}u(\eta_{i}))f^{p^{n}-1}f^{1-1/p^{e-1}}=F^{n}(F^{n}_{*}(u(\eta_{i})f^{1-1/p^{e-1+n}}))$
$=F^{n}(F^{n}_{*}\tau_{i}^{(e-1+n)}).$
By the induction hypothesis, we have $\tau_{i+1}^{(e-1)}\neq 0$, thus we have $\tau_{i}^{(e-1+n)}\neq 0$. Since $e-1+n\geq e$, we have $\tau_{i}^{(e)}\neq 0$, as desired. Next, we assume that $\tau_{i}$ is in the case (ii-i) in Definition 5.4. Then there exists an integer $n\geq 2$ such that $\Phi_{R,n}(\eta_{i})=V^{n-1}(F^{n}_{*}\eta_{i+1})$. We take regular sequence $x_{1},\ldots,x_{d}$ on $\overline{R}$ and elements $a,b\in\overline{R}$ such that $\eta_{i}=[a/x_{1}\cdots x_{d}]$, $\eta_{i+1}=[b/x_{1}^{p^{n}}\cdots x_{d}^{p^{n}}]$, and
$[a^{p}]\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{R,n}.$
Therefore, there exist elements $c_{0},\ldots,c_{n-1}\in\overline{A}$ such that
$[a^{p}]+(c_{0}f,\ldots,c_{n-1}f)\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{A,n}.$
Since we have $(c_{0}d,\ldots,c_{n-1}f)[f^{p-p^{1/p^{e+n}}}]\in[f]Q_{A_{\infty},n}$, we obtain
$[af^{1-1/p^{e+n-1}}]^{p}\equiv V^{n-1}(bf^{p^{n}-1/p^{e-1}})\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}],[f])Q_{A_{\infty},n}.$
In particular, we have
$\Phi_{A_{\infty},n}(\tau_{i}^{(e+n-1)})$ $=\left[\frac{[af^{1-1/p^{e+n-1}}]^{p}}{x_{1}\cdots x_{d}f}\right]$
$=\left[\frac{V^{n-1}(F^{n}_{*}(bf^{p^{n}-1/p^{e-1}}))}{x_{1}\cdots x_{d}f}\right]=V^{n-1}\left(F^{n}_{*}\left[\frac{bf^{p^{n}-1/p^{e-1}}}{x_{1}^{p^{n}}\cdots x_{d}^{p^{n}}f^{p^{n}}}\right]\right)$
$=V^{n-1}\left(F^{n}_{*}\left[\frac{bf^{1-1/p^{e-1}}}{x_{1}^{p^{n}}\cdots x_{d}^{p^{n}}f}\right]\right)=V^{n-1}(F^{n}_{*}\tau_{i+1}^{(e-1)}),$
which is non-zero by the induction hypothesis and $H^{d}_{\mathfrak{m}}(\overline{A}_{\infty})=0$. Therefore, we obtain $\tau_{i}^{(e)}\neq 0$ by $e+n-1\geq e$, as desired. $\blacksquare$
By Claim 5.11, we have $\tau f^{1-p^{1/p^{e}}}=\tau_{0}^{(e)}\neq 0$ for every integer $e\geq 0$. Next, let us show $R\to R_{\infty}$ is pure. By the argument of [p-pure]*Proposition 6.5, it is enough to show that $fA\to(f^{1/p^{\infty}})$ is pure. Since $(f^{1/p^{\infty}})\simeq\mathrm{colim}_{e}f^{1/p^{e}}A_{\infty}$, it is enough to show that $fA\to f^{1/p^{e}}A_{\infty}$ is pure for every $e\geq 0$. We fix an integer $e\geq 0$. Since $f$ is a regular element in $A$ and $A_{\infty}$, it is enough to show that the homomorphism $A\to A_{\infty}\xrightarrow{\cdot f^{1-1/p^{e}}}A_{\infty}$ is pure. We consider the following commutative diagram in which horizontal sequence is exact:
[…]
Since the image $\tau^{\prime}$ of $\eta$ by the map $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(\overline{A})$ is an element of the socle of $H^{d+1}_{\mathfrak{m}}(\overline{A})$, we have
$(H^{d+1}_{\mathfrak{m}}(\overline{A})\xrightarrow{\cdot f^{1-1/p^{e}}}H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty}))(\tau^{\prime})=\tau f^{1-1/p^{e}}\neq 0,$
and the image of $\tau^{\prime}$ by $H^{d+1}_{\mathfrak{m}}(\overline{A})\to H^{d+2}_{\mathfrak{m}}(A)$ is contained in the socle of $H^{d+2}_{\mathfrak{m}}(A)$, we obtain the homomorphism $H^{d+2}_{\mathfrak{m}}(A)\xrightarrow{\cdot f^{1-1/p^{e}}}H^{d+2}_{\mathfrak{m}}(A_{\infty})$ is injective. Since $A$ is Gorenstein, we obtain that the map $A\xrightarrow{\cdot f^{1-1/p^{e}}}A_{\infty}$ is pure, as desired. ∎
Theorem 5.12.
(cf. Theorem A) Let $(R,\mathfrak{m})$ be a local $\mathbb{Z}_{(p)}$-algebra such that $R$ is $p$-torsion free and $p\in\mathfrak{m}$. We assume $R$ is complete intersection. If $R$ is quasi-$F$-split, then $R$ is perfectoid pure. In particular, if $R/pR$ is quasi-$F$-split, then $R$ is perfectoid pure.
Proof.
By Proposition 4.7 and [p-pure]*Lemma 4.13, we may assume that $R$ is complete. Therefore, there exist a regular local ring $R_{0}$ and regular sequence $f_{1},\ldots,f_{r}$ such that $R_{0}/(f_{1},\ldots,f_{r})\simeq R$. Let $R_{i}:=R_{0}/(f_{1},\ldots,f_{i})$ for $r\geq i\geq 1$. We take a perfectoid $R_{0,perfd}$ as in [p-pure]*Section 4.3. By Andre’s flatness lemma [BS]*Theorem 7.14, there exists a $p$-completely faithfully flat extension $R_{0,perfd}\to R_{0,\infty}$ to a perfectoid such that $R_{0,\infty}$ has systems of $p$-power roots of $f_{1},\ldots,f_{r}$. We set $R_{i,\infty}:=(R_{0,\infty}/(f_{1},\ldots,f_{i}))_{perfd}$. Let us show that $R_{i}\to R_{i,\infty}$ is pure for each $r\geq i\geq 0$ by induction on $i$. Since $R_{0}\to R_{0,\infty}$ is faithfully flat, we obtain the case of $i=0$. We take $i\geq 1$ and assume $R_{i-1}\to R_{i-1,\infty}$ is pure. By Theorem 5.10, it is enough to show that $R_{i-1,\infty}$ is cohomologically Cohen-Macaulay and $f_{i},p$ are regular elements of $R_{i-1,\infty}$. The cohomologically Cohen-Macaulayness follows from [p-pure]*Theorem 4.24. Furthermore, the sequence $p,f$ is a regular sequence on $R_{i-1,\infty}$ by the proof of [p-pure]*Theorem 4.24, and in particular, the element $p$ is a regular element. By Lemma 5.13, the element $f$ is a regular element of $R_{i-1,\infty}$, as desired. The last assertion follows from the first assertion and Proposition 4.6. ∎
Lemma 5.13.
Let $(R,\mathfrak{m})$ be a Noetherian local ring and $M$ an $R$-module. For elements $f,g\in\mathfrak{m}$, if a sequence $f,g$ is a regular sequence on $M$ and $M$ is $(f)$-separated, then $g$ is a regular element.
Proof.
We take $a\in M$ such that $ag=0$. Since $f,g$ is a regular sequence, so is $f^{n},g$ for every integer $n\geq 1$. Since $ag=0$ in $M/f^{n}M$, we have $a\in f^{n}M$ for every integer $n\geq 1$. Since $M$ is $(f)$-separated, we have $a=0$, as desired. ∎
Lemma 5.14.
We use Notation 5.1. Let $R_{\infty}$ be a perfectoid over $R$ and $\eta\in H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$. We assume $R_{\infty}$ has a system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. If $\Phi_{R_{\infty},n}(\eta)\neq 0$ for some integer $n\geq 1$, then $p^{1-i_{n}}\eta\neq 0$, where $i_{n}=1/p+\cdots+1/p^{n}$.
Proof.
First, we consider the case where $n=1$. Then the image of $\eta$ by the composition
[…]
is non-zero. Indeed, the homomorphism $F^{-1}\colon\overline{R}_{\infty}\to R_{\infty}/(p^{1/p})$ is an isomorphism. Considering the images in $H^{d+1}_{\mathfrak{m}}(R_{\infty})$, we obtain $p^{1-1/p}(\eta/p)\neq 0$. In particular, we have $p^{1-1/p}\eta\neq 0$. Next, let us consider the case where $n\geq 2$. We may assume $\Phi_{R,n-1}(\eta)=0$. Then $p^{1-i_{n}}\eta\neq 0$ by the proof of Theorem 5.8, as desired. ∎
Theorem 5.15.
(cf. Theorems A and 6.9) We use Notation 5.1. We assume $R$ is complete intersection. Let $n$ a positive integer. Then $\operatorname{ht}(R)=n$ if and only if $R$ is perfectoid pure and
$1-\frac{1}{p}-\cdots-\frac{1}{p^{n-1}}\geq\operatorname{ppt}(R;{\rm div}(p))\geq 1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}.$
Proof.
We set $i_{m}:=1/p+\cdots+1/p^{m}$ for integers $m\geq 1$ and
$\varepsilon_{0}:=1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}=1-i_{n}+1/p^{n}(1-i_{n})+1/p^{2n}(1-i_{n})+\cdots.$
Let us show the “only if” part. Thus, we assume $\operatorname{ht}(R)=n$. First, the inequality $\operatorname{ppt}(R;{\rm div}(p))\leq i_{n-1}$ follows from Proposition 5.2, thus let us show the other inequality. If $n=1$, then $\overline{R}$ is $F$-split, thus we have $\operatorname{ppt}(R;{\rm div}(p))=1$ by Theorem 5.8, as desired. We assume $n\geq 2$ and take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. First, let us show that $\operatorname{ppt}(R;{\rm div}(p))\geq\varepsilon_{0}$. Then there exists $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(F^{n}_{*}\eta_{1})$. We take a perfectoid $R_{\infty}$ over $R$ as in the proof of Theorem 5.8. The image of $\eta_{1}$ by $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$ is denoted by $\eta^{\prime}_{1}$. If $p^{\alpha}\eta^{\prime}_{1}\neq 0$ for some $\alpha>\varepsilon_{0}$, then we have
$\operatorname{ppt}(R;{\rm div}(p))\geq 1-i_{n}+\alpha/p^{n}>\varepsilon_{0}$
by the proof of Theorem 5.8. Therefore, we may assume $\Phi_{R_{\infty},n-1}(\eta^{\prime}_{1})=0$ by Lemma 5.14. Since $R$ is perfectoid pure by Theorem 5.12, we obtain $\Phi_{R,n-1}(\eta_{1})=0$. Therefore, we obtain $\eta_{2}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta_{1})=V^{n-1}(F^{n}_{*}\eta_{2})$. Repeating such a process, we may assume that we obtain a sequence $\eta_{0}:=\eta,\eta_{1},\eta_{2},\ldots$ on $H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta_{i})=V^{n-1}(F^{n}_{*}\eta_{i+1})$ for every $i\geq 0$. We set
$\alpha_{i}:=\sup\{\alpha\in\mathbb{Z}[1/p]_{\geq 0}\mid p^{\alpha}\eta^{\prime}_{i}\neq 0\},$
where $\eta^{\prime}_{i}$ is the image of $\eta_{i}$ by $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$ for every $i\geq 0$. By the proof of Theorem 5.8, we obtain $\alpha_{i}\geq 1-i_{n}$ and $1-i_{n}+\alpha_{i+1}/p^{n}=\alpha_{i}$ for every $i\geq 0$. In other words, we obtain $\alpha_{i+1}-\varepsilon_{0}=p^{n}(\alpha_{i}-\varepsilon_{0})$. Suppose $\alpha_{0}<\varepsilon_{0}$, then for enough large $i$, we have $\alpha_{i}<1-i_{n}$, thus we have a contradiction. Therefore, we obtain $\alpha_{0}=\operatorname{ppt}(R;{\rm div}(p))\geq\varepsilon_{0}$, as desired.
Next, let us show the “if” part. Since we have
$\operatorname{ppt}(R;{\rm div}(p))\geq\varepsilon_{0}>1-i_{n},$
we obtain $\operatorname{ht}(R)\leq n$ by Proposition 5.2. Furthermore, if $n\geq 2$ and $\operatorname{ht}(R)\leq n-1$, then we obtain a contradiction by the “only if” part. Therefore, we obtain $\operatorname{ht}(R)=n$, as desired. ∎
Corollary 5.16.
(cf. Theorem A) We use Notation 5.1. We assume $R$ is complete intersection. Let $n$ a positive integer. Then $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$ if and only if $R$ is quasi-$(F,F^{\infty})$-split of height $n$.
Proof.
The “if” part follows from Theorem 5.8. We assume $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, then $\operatorname{ht}(R)=n$ by Theorem 5.15. We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Then there exists $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(F^{n}_{*}\eta_{1})$. We take a perfectoid $R_{\infty}$ over $R$ as in the proof of Theorem 5.8. Since $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, we have $p^{\alpha}\eta_{1}^{\prime}\neq 0$ for $1>\alpha\geq 0$ by the proof of Theorem 5.8, where $\eta^{\prime}_{1}$ is the image of $\eta_{1}$ by $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$. For an integer $e\geq 1$, since the composition
$H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\xrightarrow{F^{e}}H^{d}_{\mathfrak{m}}(F^{e}_{*}\overline{R}_{\infty})\xrightarrow{F^{-e}}H^{d}_{\mathfrak{m}}(R_{\infty}/(p^{1/p^{e}}))$
coincides with the natural homomorphism $H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\to H^{d}_{\mathfrak{m}}(R_{\infty}/(p^{1/p^{e}}))$ and $F^{-e}\colon\overline{R}_{\infty}\to R_{\infty}/(p^{1/p^{e}})$ is an isomorphism, we obtain $F^{e}(\eta_{1}^{\prime})\neq 0$. Therefore, we obtain $\Phi_{R,n,e}(\eta)\neq 0$ for every $e\geq 1$, thus $R$ is quasi-$(F,F^{\infty})$-split of height $n$, as desired. ∎
Corollary 5.17.
(cf. [p-pure]*Proposition 6.7) Let $(R,\mathfrak{m})$ be a Noetherian complete local domain with $p\in\mathfrak{m}\backslash\{0\}$. We assume $R$ is complete intersection. If $R/pR$ is quasi-$F$-split and $R[1/p]$ is regular, then $R$ is BCM-regular.
Proof.
If $R$ is not BCM-regular, then $(R,p^{\varepsilon}{\rm div}(p))$ is not perfectoid pure by the proof of [p-pure]*Proposition 6.7 for every $\varepsilon\in\mathbb{Z}[1/p]_{>0}$. Therefore, $R$ is BCM-regular by Theorem 5.15. ∎
6. Case of graded rings
In this section, we study quasi-$F$-splitting of graded rings. The results vary greatly depending on an invariant called a-invariant.
Notation 6.1.
Let $S=\bigoplus_{i\in\mathbb{Q}_{\geq 0}}S_{i}$ be a Noetherian graded ring such that $S_{0}$ is a divisorial valuation ring with $p$ in the maximal ideal of $S_{0}$, $S$ is $p$-torsion free, and $S_{i}=0$ if $i\notin\mathbb{Z}$. We set $\mathfrak{m}:=(p,S_{>0})$, $R:=S_{\mathfrak{m}}$, and $d:=\dim{\overline{R}}$. Then $R$ satisfies the conditions in Notation 5.1. Furthermore, we define $S^{(m)}:=\bigoplus_{i\geq 0}S_{im}$ as graded ring for every $m\in\mathbb{Q}_{>0}$. For a graded $S$-module $M$ and $m\in\mathbb{Q}$, we define a graded $S$-module $M(m)$ by $M(m)_{i}=M_{i+m}$ for all $i\in\mathbb{Q}$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\simeq H^{d}_{\mathfrak{m}}(\overline{S})$, we endow $H^{d}_{\mathfrak{m}}(\overline{R})$ with a graded structure. We set $a(\overline{S}):=\max\{m\in\mathbb{Z}\mid H^{d}_{\mathfrak{m}}(\overline{S})_{m}\neq 0\}$. Furthermore, we define the $\mathbb{Q}$-graded structure on $W_{n}(S)$ by
$W_{n}(S)_{i}:=\{(a_{0},\ldots,a_{n-1})\in W_{n}(S)\mid a_{0}\in S_{i},\ldots,a_{n-1}\in S_{p^{n-1}i}\}$
for $i\in\mathbb{Q}$ (cf. [KTTWYY2]*Proposition 7.1). We consider the graded structure on $Q_{S,n}$ by $\overline{W}_{n}(S)^{(p)}$. Then we also endow $H^{d}_{\mathfrak{m}}(Q_{R,n})$ with a graded structure.
6.1. Case of $a(\overline{S})<0$
In this subsection, we study the case where $a(\overline{S})<0$, and the case is related to the cone of log Fano pairs (cf. Corollaries 6.4 and 6.5).
Proposition 6.2.
We use Notation 6.1. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ be a homogeneous element of degree $\nu$. Then we can run a program in Definition 5.4 such that if we obtain $\eta_{1},\ldots,\eta_{i}$, then there exist $0=e_{0}<e_{1}<\cdots<e_{i}$ such that each $\eta_{j}\in H^{d}_{\mathfrak{m}}(\overline{R})$ is homogeneous of degree $p^{e_{j}}\nu$ for $i\geq j\geq 0$.
Proof.
We note that $\Phi_{S,n}\colon S\to Q_{S,n}$, $V^{n-1}\colon F^{n}_{*}\overline{S}^{(p^{n})}\to Q_{S,n}$ and $F^{e}\colon\overline{S}\to F^{e}_{*}\overline{S}^{(p^{e})}$ are homomorphisms of graded $S$-modules. We assume that we get a homogeneous element $\eta_{j-1}$ of degree $p^{e_{j-1}}\nu$ by a program in Definition 5.4 for an integer $j\geq 1$. If $\eta_{j-1}$ is in the case (i-i), then $\eta_{j}=F^{e}(\eta_{j})$ for some $e\geq 1$, then $\eta_{j}$ is homogeneous of degree $p^{e_{j-1}+e}\nu$, thus $e_{j}:=e_{j-1}+e$ satisfies the desired condition. If $\eta_{j-1}$ is in the case (ii-i), then there exist an integer $n\geq 2$ and a homogeneous element $\eta^{\prime}$ such that $\Phi_{R,n}(\eta_{j-1})=V^{n-1}(F^{n}_{*}\eta^{\prime})$. Then $\eta^{\prime}$ is of degree $p^{e_{j-1}+n}\nu$, thus $e_{j}:=e_{j-1}+n$ and $\eta_{j}:=\eta^{\prime}$ satisfy the desired condition. ∎
Theorem 6.3.
(Theorem B) We use Notation 6.1. We assume $\overline{S}$ is normal quasi-Gorenstein and satisfies $a(\overline{S})<0$. We further assume that $\operatorname{Spec}{\overline{S}}\backslash\{\mathfrak{m}\overline{S}\}$ is strongly $F$-regular. If $R$ is quasi-$F$-split, then $R$ is perfectoid BCM-regular. Furthermore, if $p=2$, the converse implication holds.
Proof.
We take a perfectoid BCM-algebra $B$ over $R$. First, let us show that $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective. We take a homogeneous element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Since $a(\overline{S})<0$, the degree of $\eta$ is negative. It is enough to show that $\eta$ is not contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. Suppose $\eta$ is contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. We run a program as in Proposition 6.2 and get $\eta_{0}:=\eta,\eta_{1},\ldots,\eta_{i}$. Since $R$ is quasi-$F$-split, $\eta_{i}$ is not in the case (ii-ii). By Lemma 5.6, the element $\eta_{i}$ is contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. In particular, we have $\eta_{i}\in 0^{*}_{H^{d}_{\mathfrak{m}}(\overline{R})}$ by the proof of [bcm-reg]*Proposition 5.3. By Theorem B.1, there exists a positive integer $e^{\prime}$ such that $F^{e^{\prime}}(\eta_{i})=0$, and in particular, $\eta_{i}$ is not in the case (i-ii). Therefore, by a program, we obtain a sequence of homogeneous elements $\{\eta_{i}\mid i\geq 0\}$ such that there exist $e_{o}=0<e_{1}<\ldots$ with $\deg(\eta_{i})=p^{e_{i}}\deg(\eta)$ for every $i\geq 0$. Since $\overline{R}$ is $F$-pure outside of $\mathfrak{m}\overline{R}$, the kernel of $F\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(F_{*}\overline{R})$ is of finite length, and in particular, there exists $j\geq 0$ such that $F^{e}(\eta_{j})\neq 0$ for every $e\geq 1$. However, it contradicts the fact that the program does not stop. In conclusion, we obtain the injectivity of $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$. Since $\overline{S}$ is quasi-Gorenstein, so is $\overline{R}$, and in particular, the homomorphism $\overline{R}\to\overline{B}$ is pure. Since $\overline{B}$ is a BCM-algebra over $\overline{R}$, we obtain that $\overline{R}$ is Gorenstein, thus so is $R$. Since $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{B})$ is injective and $B$ is a BCM-algebra over $R$, we obtain the injectivity of $H^{d+1}_{\mathfrak{m}}(R)\to H^{d+1}_{\mathfrak{m}}(B)$, thus the homomorphism $R\to B$ is pure, as desired. Next, we assume $p=2$ and $R$ is BCM-regular. Since $R$ is Gorenstein, $R$ is quasi-$F$-split by Corollary 5.3. ∎
Corollary 6.4.
Let $k$ be an algebraically closed field of characteristic $p$. Let $X$ be a normal integral projective flat scheme over $\operatorname{Spec}{W(k)}$. We assume $-K_{X}$ is ample $\mathbb{Q}$-Cartier. We set
$S:=\bigoplus_{m\geq 0}H^{0}(X,\mathcal{O}_{X}(-mK_{X})).$
The closed fiber of $X\to\operatorname{Spec}{W(k)}$ is denoted by $\overline{X}$ and we assume that $\overline{X}$ is klt. We assume one of the following conditions:
  1. (1)
    $\overline{X}$ is strongly $F$-regular and quasi-$F$-split, or
  2. (2)
    $\dim{\overline{X}}=2$ and $p>5$.
Then $S_{(p,S_{>0})}$ is perfectoid BCM-regular.
Proof.
By [KTTWYY2]*Theorem 6.3 and [Hara2]*Theorem 1.1, we may assume $\overline{X}$ is strongly $F$-regular and quasi-$F$-split. We set $d:=\dim(X)$ and
$S^{\prime}:=\bigoplus_{m\geq 0}H^{0}(\overline{X},\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})).$
Then $S^{\prime}$ is normal, quasi-Gorenstein by [watanabe81]*Theorem 2.8, and quasi-$F$-split by [KTTWYY2]*Theorem 7.16 and [kty]*Proposition 2.25. Since we have
$H^{d}_{\mathfrak{m}}(S^{\prime})_{m}\simeq H^{d-1}(\overline{X},\mathcal{O}_{\overline{X}}(-mK_{\overline{X}}))\simeq H^{0}(\overline{X},\mathcal{O}_{\overline{X}}((m+1)K_{\overline{X}}))$
and $-K_{\overline{X}}$ is ample, we obtain $a(\overline{S^{\prime}})<0$. Furthermore, since $\overline{X}$ is strongly $F$-regular, so is $\operatorname{Spec}{S^{\prime}}\backslash\{S^{\prime}_{>0}\}$. Therefore, it is enough to show that $S/pS\simeq S^{\prime}$ as graded rings by Theorem 6.3. We note that if $S/pS\simeq S^{\prime}$, then we have
$H^{0}(X,\mathcal{O}_{X})\otimes_{W(k)}k\simeq S_{0}/pS_{0}\simeq S^{\prime}_{0}\simeq H^{0}(\overline{X},\mathcal{O}_{\overline{X}})\simeq k,$
and in particular, $S_{0}$ is divisorial valuation ring with the maximal ideal $(p)$. Furthermore, by [bcm-reg]*Theorem 6.27, every stalk of $\mathcal{O}_{X}$ is perfectoid BCM-regular, and in particular, the sheaf $\mathcal{O}_{X}(-mK_{X})$ is Cohen-Macaulay. Since we have $\mathcal{O}_{X}(-mK_{X})\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})$ on the codimension one point of $\overline{X}$ and $\mathcal{O}_{X}(-mK_{X})$ satisfies the condition $S_{3}$ for each integer $m$, we obtain $\mathcal{O}_{X}(-mK_{X})\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})$. Since $\overline{X}$ is quasi-$F$-split, we obtain $H^{1}(\overline{X},\mathcal{O}_{\overline{X}}(-mK_{\overline{X}}))=0$ for every integer $m\geq 0$ by [KTTWYY1]*Theorem 3.15. By [hartshorne]*Theorem 12.11, we obtain
$H^{0}(X,\mathcal{O}_{X}(-mK_{X}))\otimes_{W(k)}k\simeq H^{0}(\overline{X},\mathcal{O}_{\overline{X}}(-mK_{\overline{X}})),$
thus, we obtain $S/pS\simeq S^{\prime}$, as desired. ∎
Corollary 6.5.
Let $k$ be an algebraically closed field of characteristic $p$. Let $X$ be a normal integral projective flat scheme over $\operatorname{Spec}{W(k)}$ and $\Delta$ an effective $\mathbb{Q}$-Weil divisor on $X$, whose components are flat over $\operatorname{Spec}{W(k)}$. We assume that $-(K_{X}+\Delta)$ is ample $\mathbb{Q}$-Cartier and $\Delta$ has standard coefficients. We set
$S:=\bigoplus_{m\geq 0}H^{0}(X,\mathcal{O}_{X}(-m(K_{X}+\Delta))).$
The closed fiber of $(X,\Delta)\to\operatorname{Spec}{W(k)}$ is denoted by $(\overline{X},\overline{\Delta})$ and we assume that $(\overline{X},\overline{\Delta})$ is strongly $F$-regular, the Picard rank of $\overline{X}$ is one, and $\overline{X}$ is $\mathbb{Q}$-factorial. We assume one of the following conditions:
  1. (1)
    $(\overline{X},\overline{\Delta})$ is quasi-$F$-split,
  2. (2)
    $\dim{\overline{X}}=1$, or
  3. (3)
    $\dim{\overline{X}}=2$ and $p>41$.
Then $S_{(p,S_{>0})}$ is perfectoid BCM-regular.
Proof.
By [KTTWYY1]*Corollary 5.16 and [KTTWYY2]*Theorem C, we have $(\overline{X},\overline{\Delta})$ is quasi-$F$-split. By the same argument as the proof of Corollary 6.4, it is enough to show that for every integer $m\geq 0$,
  1. (a)
    $\mathcal{O}_{X}(-m(K_{X}+\Delta))\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta}))$,
  2. (b)
    $H^{1}(\overline{X},\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta})))=0$, and
  3. (c)
    $H^{d-1}(\overline{X},\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta})))=0$, where $d:=\dim{X}$.
Indeed, the cone of $(\overline{X},\overline{\Delta})$ coincides with $\overline{S}$ by (a) and (b), the $a$-invariant $a(\overline{S})$ is negative by (c). By an argument in the proof of Corollary 6.4, we have that $X$ is divisorially Cohen-Macaulay, that is, for all divisorial sheaves on $X$ are Cohen-Macaulay. We set $\Delta:=\sum a_{i}E_{i}$, where each $E_{i}$ is a prime divisor and $a_{i}$ has a form $a_{i}=(m_{i}-1)/m_{i}$ for some integer $m_{i}\geq 2$. Since $(\overline{X},\overline{\Delta})$ is strongly $F$-regular, we have $\lfloor\overline{\Delta}\rfloor=0$, and in particular, we have $E_{i}|_{\overline{X}}$ is a reduced divisor. Therefore, we obtain that $\mathcal{O}_{X}(-m(K_{X}+\Delta))\otimes_{\mathcal{O}_{X}}\mathcal{O}_{\overline{X}}\simeq\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta}))$ on codimension one points of $\overline{X}$. Since $X$ is divisorially Cohen-Macaulay, we obtain the condition (a). Next, let us show the condition (b). We take an integer $m\geq 0$ and set $L:=\lfloor-m(K_{X}+\Delta)\rfloor$, which is $\mathbb{Q}$-Cartier since $\overline{X}$ is $\mathbb{Q}$-factorial. Since the Picard rank of $\overline{X}$ is one and $-K_{X}$ is ample, we obtain that $L-K_{X}$ or $-L$ is ample. By [KTTWYY1]*Theorem 5.1, we obtain the condition (b) if $\dim{\overline{X}}\geq 2$. Therefore, it is enough to show the condition (c). By the Serre duality, we obtain
$H^{d-1}(\overline{X},\mathcal{O}_{\overline{X}}(-m(K_{\overline{X}}+\overline{\Delta})))\simeq H^{0}(\overline{X},\mathcal{O}_{\overline{X}}((m+1)K_{\overline{X}}+\lceil m\overline{\Delta}\rceil)).$
Since $-(K_{\overline{X}}+\overline{\Delta})$ is ample, it is enough to show that $\lceil m\overline{\Delta}\rceil\leq(m+1)\overline{\Delta}$ for every integer $m\geq 0$. We take a prime divisor $E$ on $\overline{X}$ and a positive integer $n$ with $\frac{n-1}{n}=\mathrm{ord}_{E}(\overline{\Delta})$. We take integers $q\geq 0$ and $n>r\geq 0$ such that $m=nq+r$. Then we have
$\lceil m\frac{n-1}{n}\rceil=m-\lfloor\frac{m}{n}\rfloor=m-q.$
On the other hand, we have
$(m+1)\frac{n-1}{n}=m+1-q-\frac{1}{n},$
thus we obtain $(m+1)\frac{n-1}{n}\geq\lceil m\frac{n-1}{n}\rceil$, as desired. ∎
Corollary 6.6.
(Theorem C) There exists a Noetherian graded ring $S=\bigoplus_{i\in\mathbb{Z}_{\geq 0}}S_{i}$ such that $S$ satisfies the following properties.
  1. (1)
    $S$ is torsion free and $S_{0}$ is a divisorially valuation ring with maximal ideal $(p)$.
  2. (2)
    We set $\mathfrak{m}:=(S_{>0},p)$ and $R:=S_{\mathfrak{m}}$, then $R$ is perfectoid BCM-regular.
  3. (3)
    $R/pR$ is normal Gorenstein but not strongly $F$-regular.
Proof.
We set $k:=\overline{\mathbb{F}}_{p}$. If $p=2,3$ or $5$, then we consider the pair
$(\mathbb{P}^{1}_{W(k)},\Delta:=\frac{1}{2}E_{1}+\frac{2}{3}E_{2}+\frac{4}{5}E_{3}),$
where $E_{1},E_{2},E_{3}$ are disjoint sections of $\mathbb{P}^{1}_{W(k)}\to\operatorname{Spec}{W(k)}$. Then it satisfies the assumptions of Corollary 6.5, thus the localization of
$S:=\bigoplus_{m\geq 0}H^{0}(X,\mathcal{O}_{X}(-mK_{X}))$
is perfectoid BCM-regular. On the other hand, we see that $S=W(k)[x,y,z]/(z^{2}+x^{3}+y^{5})$, thus $S/pS$ is not strongly $F$-regular by [Hara2]*Theorem 1.1. Therefore, the ring $S$ satisfies the desired conditions. Next, we assume $p\geq 7$. By [CTW18]*Theorem 1.1, there exists a projective klt surface $\overline{X}$ such that $-K_{\overline{X}}$ is ample and $\overline{X}$ is not globally $F$-regular. By [CTW18]*Theorem 1.1 and [ABL]*Proposition 2.5, a log resolution of $\overline{X}$ lifts to $W(k)$. By [BBKW]*Proposition 6.2, there exists a flat normal integral projective scheme $X$ over $\operatorname{Spec}{W(k)}$ such that $X\times_{\operatorname{Spec}{W(k)}}\operatorname{Spec}{k}\simeq\overline{X}$ and $K_{X}$ is $\mathbb{Q}$-Cartier. Since $-K_{\overline{X}}$ is ample, so is $-K_{X}$. By Corollary 6.4, the localization $R$ of
$S:=\bigoplus_{m\geq 0}H^{0}(X,\mathcal{O}_{X}(-mK_{X}))$
is perfectoid BCM-regular. Furthermore, by the proof of Corollary 6.4, $S/pS$ is cone of $\overline{X}$, thus $S/pS$ is not strongly $F$-regular by [smith]*Theorem 3.10. Thus, $R$ is perfectoid BCM-regular, but $R/pR$ is not strongly $F$-regular. Moreover, the ring $S_{0}$ is the divisorially valuation ring with maximal ideal $(p)$, thus $S$ satisfies the desired conditions. ∎
6.2. Case of $a(\overline{S})\geq 0$
In this subsection, we study the case where $a(\overline{S})\geq 0$.
Proposition 6.7.
We use Notation 6.1. If $R$ is quasi-$F$-split, then we have $a(R)\leq 0$.
Proof.
Suppose $a(R)>0$, there exists a homogeneous non-zero element $\widetilde{\eta}\in H^{d+1}_{\mathfrak{m}}(R)$ of positive degree. Replacing $\widetilde{\eta}$ by $p^{s}\widetilde{\eta}$ for some $s\geq 0$, we may assume $\widetilde{\eta}$ is $p$-torsion free. Then there exists a homogeneous element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of positive degree such that $\widetilde{\eta}=\eta/p$. We assume that we get $\eta_{0}:=\eta,\eta_{1},\ldots,\eta_{i}$ by a program as in Proposition 6.2, then $\eta_{j}$ is homogeneous of positive degree for every $i\geq j\geq 0$. Since $R$ is quasi-$F$-split, $\eta_{i}$ is not in the case (ii-ii). Since the degree of $H^{d}_{\mathfrak{m}}(\overline{R})$ is upper bounded, there exists a positive integer $e$ such that $F^{e}(\eta_{i})=0$, thus $\eta_{i}$ is not in case (i-ii). In conclusion, we get $\{\eta_{i}\mid i\geq 0\}$ by such a program. However, by Proposition 6.2, the degrees of $\eta_{i}$ are strictly increasing, thus it contradicts $\eta_{i}\neq 0$ for every $i\geq 0$, as desired. ∎
Proposition 6.8.
We use Notation 6.1. We assume $R$ is Gorenstein and $a(\overline{S})=0$. Then
  1. (1)
    $\operatorname{ht}(R)=\operatorname{ht}(\overline{R})$ and
  2. (2)
    $R$ is quasi-$(F,F^{\infty})$-split if and only if $\overline{R}$ is $F$-pure.
Proof.
Let us show assertion (1). By Proposition 4.6, we have $\operatorname{ht}(R)\leq\operatorname{ht}(\overline{R})$. We set $n:=\operatorname{ht}(R)$, then it is enough to show that $\overline{R}$ is $n$-quasi-$F$-split. If $n=1$, the assertion follows from $Q_{R,1}\simeq Q_{\overline{R},1}\simeq F_{*}\overline{R}$, thus we assume $n\geq 2$. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous non-zero element of degree zero, then $\eta$ is contained in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ by $a(\overline{S})=0$. We consider the following commutative diagram;
[…]
Since $\Phi_{R,n}(\eta)\neq 0$, $\Phi_{R,n-1}(\eta)=0$, and $R$ is Cohen-Macaulay, there uniquely exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(\eta^{\prime})$. By Proposition 6.2, the degree of $\eta^{\prime}$ is zero. By the exact sequence
$0\to F^{n-1}_{*}B_{\overline{R}}\to Q_{\overline{R},n}\to F_{*}\overline{R}\to 0$
and the Cohen-Macaulayness of $\overline{R}$, the homomorphism $V^{n-1}\colon H^{d}_{\mathfrak{m}}(F^{n-1}_{*}B_{\overline{R}})\to H^{d}_{\mathfrak{m}}(Q_{\overline{R},n})$ is injective, thus it is enough to show that $\eta^{\prime}$ is not contained in the kernel of $\alpha$. By the exact sequence
$0\to F^{n-1}_{*}\overline{R}\xrightarrow{F^{n-1}_{*}F}F^{n}_{*}\overline{R}\to F^{n-1}_{*}B_{\overline{R}}\to 0,$
[…]
Next, let us show assertion (2). If $\overline{R}$ is $F$-pure, then $R$ is quasi-$(F,F^{\infty})$-split by definition. We assume $R$ is quasi-$(F,F^{\infty})$-split and set $n:=\operatorname{ht}(R)$. Suppose $n\geq 2$. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a homogeneous non-zero element of degree zero, then it is contained in the socle. We consider the following diagram;
[…]
By the argument in the proof of (1), there exists a homogeneous element $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(\overline{R})$ of degree zero such that $\Phi_{R,n}(\eta)=V^{n-1}(\eta^{\prime})$. Since $\overline{R}$ is not $F$-pure, if $e\geq 2$, then $F^{e-1}(\eta^{\prime})=0$, and it contradicts the fact that $R$ is quasi-$(F,F^{\infty})$-split. Therefore, we obtain that $n=1$, as desired. ∎
Theorem 6.9.
(Theorem D) We use Notation Notation 6.1. We assume $R$ is complete intersection and $a(\overline{S})=0$. If $\operatorname{ht}(\overline{R})=n<\infty$, then $R$ is perfectoid pure and we have
$\operatorname{ppt}(R;{\rm div}(p))=1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}.$
Proof.
By Proposition 6.8, we obtain $\operatorname{ht}(R)=n$. By Theorem 5.12, the ring $R$ is perfectoid pure. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a non-zero element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$, then $\deg(\eta)=0$. We consider a program as in Proposition 6.2 and get $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$, then $\deg(\eta_{1})=0$, and in particular, there exists $u\in\overline{R}^{\times}$ such that $\eta_{1}=u\eta$. If $\eta$ is in the case (i), then $\overline{R}$ is $F$-pure, thus the assertion follows from Theorem 5.8 and Proposition 6.8. We assume $\eta$ is in the case (ii). Since $\operatorname{ht}(R)=n$, we have $\Phi_{R,n}(\eta)=V^{n-1}(F^{n}_{*}\eta_{1})$. We take a perfectoid $R_{\infty}$ as in the proof of Theorem 5.8. We take a system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$ in $R_{\infty}$ and $T\in R_{\infty}^{\flat}$ is the corresponding element of the projective system $\{p^{1/p^{e}}\}_{e\geq 0}$. We set $\tau:=(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))(\eta)$ and $\eta^{\prime}_{1}:=(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))(\tau_{1})$. By the proof of Theorem 5.8, we have
(6.1) $p^{(1-i_{n})}(\tau/p)=wF^{-n}(\tau)/(p^{1/p^{n}}),$
where $w$ is an unit of $R_{\infty}$ and $i_{n}:=1/p+\cdots+1/p^{n}$. We set $\alpha:=\operatorname{ppt}(R;{\rm div}(p))\geq 0$. We consider the diagram
[…]
where $\mathfrak{m}^{\flat}$ is the ideal of $R_{\infty}^{\flat}$ containing $T$ and $\mathfrak{m}^{\flat}/(T)\simeq\mathfrak{m}/(p)$, then the corresponding element to $\tau$ is denoted by $\tau^{\flat}\in H^{d+1}_{\mathfrak{m}}(R_{\infty}^{\flat})$. By (6.1), we have $T^{(1-i_{n})}\tau^{\flat}=w^{\prime}F^{-n}(\tau^{\flat})$ for some unit $w^{\prime}\in R_{\infty}^{\flat}$. Since $F^{-n}$ is an isomorphism, we have $T^{\alpha}\tau^{\flat}=0$ if and only if $T^{1-i_{n}+\alpha/p^{n}}\tau^{\flat}=0$ for $\alpha\in\mathbb{Z}[1/p]_{\geq 0}$. We set $\varepsilon:=\operatorname{ppt}(R;{\rm div}(p))$, then for $\alpha\in\mathbb{Z}[1/p]_{\geq 0}$, we have
$\begin{cases}T^{\alpha}\tau^{\flat}=0&\alpha>\operatorname{ppt}(R;{\rm div}(p))\\ T^{\alpha}\tau^{\flat}\neq 0&\alpha<\operatorname{ppt}(R;{\rm div}(p)).\end{cases}$
Suppose $\operatorname{ppt}(R;{\rm div}(p))>1-(p+\cdots+p^{n-1})/(p^{n}-1)=:\varepsilon_{0}$, then there exists $\alpha\in\mathbb{Z}[1/p]$ such that $\operatorname{ppt}(R;{\rm div}(p))>\alpha>\varepsilon_{0}$. We define a sequence $\{\alpha_{m}\}_{m\geq 1}$ by $\alpha_{1}=\alpha$ and
$\alpha_{m+1}=p^{n}(\alpha_{m}-1+i_{n}),$
then we have $1-i_{n}+\alpha_{m+1}/p^{n}=\alpha_{m}$, and in particular, we have $T^{\alpha_{m}}\tau^{\flat}=0$ for every $m\geq 1$. Since we have $(1-1/p^{n})\varepsilon_{0}=1-i_{n}$, we obtain $\alpha_{m+1}-\varepsilon_{0}=p^{n}(\alpha_{m}-\varepsilon_{0})$. Since $\alpha>\varepsilon_{0}$, there exists $m\geq 1$ such that $\alpha_{m}>\operatorname{ppt}(R;{\rm div}(p))$, thus we get a contradiction. Next, we suppose $\operatorname{ppt}(R;{\rm div}(p))<\varepsilon_{0}$, then by the same argument as above, we obtain a contradiction. Therefore, we have $\operatorname{ppt}(R;{\rm div}(p))=\varepsilon_{0}$, as desired. ∎
Proposition 6.10.
Let $R:=\mathbb{Z}_{p}[[x,y,z]]/(x^{3}+y^{3}+z^{3})$ and $p\equiv 2\ \textrm{mod}\ 3$.
  1. (1)
    The ring $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1-p/(p^{2}-1)$.
  2. (2)
    The ring $R\otimes_{\mathbb{Z}_{p}}R$ is not perfectoid pure.
Proof.
We note that $\operatorname{ht}(\overline{R})=2$. The first assertion follows from Theorem 6.9. We use same notations $R_{\infty}$, $\eta\in H^{2}_{\mathfrak{m}}(\overline{R})$, $\tau\in H^{2}_{\mathfrak{m}}(\overline{R}_{\infty})$, and $\tau^{\flat}\in H^{3}_{\mathfrak{m}^{\flat}}(R^{\flat}_{\infty})$ in the proof of Theorem 6.9. By the proof of Theorem 6.9, we obtain
(6.2) $T^{(1-i_{n})}\tau^{\flat}=w^{\prime}F^{-2}(\tau^{\flat}).$
Suppose $R\otimes_{\mathbb{Z}_{(p)}}R$ is perfectoid pure, then so is $R^{\prime}:=(R\otimes_{\mathbb{Z}_{p}}R)^{\wedge\mathfrak{n}}$ by [p-pure]*Lemma 4.8, where $\mathfrak{n}$ be the ideal of $R\otimes_{\mathbb{Z}_{p}}R$ generated by $p$, $\mathfrak{m}\otimes R$, and $R\otimes\mathfrak{m}$. We note that
$R^{\prime}\simeq\mathbb{Z}_{p}[[x,y,z,x^{\prime},y^{\prime},z^{\prime}]]/(x^{3}+y^{3}+z^{3},x^{\prime 3}+y^{\prime 3}+z^{\prime 3}).$
We take a surjection $S^{\prime}:=\mathbb{Z}_{p}[[x,y,z,x^{\prime},y^{\prime},z^{\prime}]]\to R^{\prime}$ and we define perfectoids $S^{\prime}_{\infty}$ and $R^{\prime}_{\infty}:=R^{S^{\prime}_{\infty}}_{perfd}$ as in [p-pure]*Section 4.3. By the natural homomorphism $H^{3}_{\mathfrak{m}}(R)\otimes_{\mathbb{Z}_{p}}H^{3}_{\mathfrak{m}}(R)\to H^{5}_{\mathfrak{n}}(R^{\prime})$, the image of the element $[z^{2}/xyp]\otimes[z^{\prime 2}/x^{\prime}y^{\prime}p]$ is
$[z^{2}z^{\prime 2}/xyx^{\prime}y^{\prime}p],$
which a non-zero element of the socle of $H^{5}_{\mathfrak{n}}(R^{\prime})$. Since $R^{\prime}$ is perfectoid pure, the image $\eta^{\prime}$ in $H^{5}_{\mathfrak{m}}(R^{\prime})$ is non-zero. We set
$\tau^{\prime}:=(H^{5}_{\mathfrak{n}}(R^{\prime})\to H^{5}_{\mathfrak{n}}(R^{\prime}_{\infty}))(\eta^{\prime})$
and denote $\tau^{\prime\flat}\in H^{5}_{\mathfrak{n}}((R^{\prime})_{\infty}^{\flat})$ corresponding to $\tau^{\prime}$. Then we have
$T^{2(1-i_{n})}\tau^{\prime\flat}=w^{\prime\prime}F^{-2}(\tau^{\prime\flat})$
for some unit $w^{\prime\prime}\in(R^{\prime})_{\infty}^{\flat}$ by (6.2). In particular, we have $T^{1/p}\tau^{\prime\flat}\neq 0$, thus we have $\operatorname{ppt}(R^{\prime};{\rm div}(p))\geq 1/2$. On the other hand, since $R^{\prime}$ is not quasi-$F$-split by Example 4.14, thus $\operatorname{ppt}(R;{\rm div}(p))=0$ by Proposition 5.2. Therefore, we obtain a contradiction. ∎
Appendix A Construction of functorial test perfectoids
In this section, we prove the existence of the functorial test perfectoid (Theorem A.1).
Theorem A.1.
There exists a functor $T$ from the category of $\mathbb{Z}_{(p)}$-algebras to the category of perfectoids such that $T(R)$ is a test perfectoid over $R$ and $T(R)$ has a compatible system of $p$-power roots. If $R$ is $p$-torsion free, then so is $T(R)$.
Proof.
First, we construct a perfectoid $T(R)$ over a $\mathbb{Z}_{(p)}$-algebra $R$. We set $A:=\mathbb{Z}_{p}[X]^{\wedge(p,d)}$ and $d:=X-p$, and we define the $\mathbb{Z}_{p}$-algebra homomorphism $\phi\colon A\to A$ by $\phi(X)=X^{p}$, then $(A,d)$ is a prism. We define $A_{\infty}$ by the $(p,d)$-adic completion of perfection of $A$. Let $A\{R\}:=A\{X_{f}\mid f\in R\}$ be a free $\delta$-ring as in [BS]*Lemma 2.11, then $(A\{R\}^{\wedge(p,d)},d)$ is a prim. Taking a perfection and $(p,d)$-completion, we obtain a perfect prism $(A\{R\}_{\infty},(d))$. We set $Z\{R\}_{\infty}:=A\{R\}_{\infty}/(d)$, then we have the $\mathbb{Z}_{(p)}$-algebra homomorphism $\mathbb{Z}_{(p)}[R]:=\mathbb{Z}_{(p)}[X_{f}\mid f\in R]\to A\{R\}_{\infty}$ defined by $X_{f}\mapsto X_{f}$. We set
$T(R):=(Z\{R\}_{\infty}\otimes_{\mathbb{Z}_{(p)}[R]}R)_{perfd},$
where $\mathbb{Z}_{(p)}[R]\to R$ is defined by $X_{f}\mapsto f$. We note that $Z\{R\}_{\infty}$ is a perfectoid and $\mathbb{Z}_{(p)}[R]\to R$ is surjective, the perfectoidazation exists by [BS]*Corollary 7.3. Furthermore, since the images of $\{\phi^{-e}(X)\}$ by the map $A\{R\}_{\infty}\to T(R)$ form a compatible system of $p$-power roots of $p$.
Next, let us show that $T(R)$ is a test perfectoid over $R$. Let $B$ be a perfectoid $R$-algebra, then by Andre’s flatness lemma [BS]*Theorem 7.14, we may assume that $B$ contains a compatible system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. Let $(\widetilde{B},(d^{\prime}))$ is a perfect prism such that $\widetilde{B}/(d^{\prime})\simeq B$. We consider the map $A_{\infty}=\mathbb{Z}_{(p)}[X^{1/p^{\infty}}]^{\wedge(p,d)}\to B$ by $X^{1/p^{e}}\mapsto p^{1/p^{e}}$, then it induces the homomorphism of perfectoids $A_{\infty}/(d)\to B$, and in particular, we obtain the homomorphism of perfect prisms $(A_{\infty},(d))\to(\widetilde{B},(d^{\prime}))$ by [BS]*Theorem 3.10. For each element $f\in R$, we take a lift $\widetilde{f}\in\widetilde{B}$ of the image of $f$ by $R\to B$, then we can define the homomorphism of perfect prisms $(A\{R\}_{\infty},(d))\to(\widetilde{B},(d^{\prime}))$ defined by $X_{f}\mapsto\widetilde{f}$, thus we obtain the homomorphism $Z\{R\}_{\infty}\to B$ and the commutative diagram
[…]
In conclusion, we obtain an $R$-algebra homomorphism $T(R)\to B$, as desired.
Next, for a $\mathbb{Z}_{(p)}$-algebra homomorphism $\varphi\colon R\to S$, we define a ring homomorphism $T(\varphi)\colon T(R)\to T(S)$. We define $Z\{R\}_{\infty}$ and $Z\{S\}_{\infty}$ as above, then we obtain the $\mathbb{Z}_{(p)}$-algebra homomorphism $Z\{R\}_{\infty}\to Z\{S\}_{\infty}$ by $X_{f}\mapsto X_{\varphi(f)}$. Then it induces the homomorphism $T(\varphi)\colon T(R)\to T(S)$. Then this $T$ defines the functor from the category of $\mathbb{Z}_{(p)}$-algebras to the category of perfectoids, as desired.
Finally, let us show that $T(R)$ is $p$-torsion free if $R$ is $p$-torsion free. Since the endomorphism $\phi\colon A\{R\}\to A\{R\}$ is $(p,d)$-completely faithfully flat, so is $A\{R\}\to A\{R\}_{\infty}$. Since $A[R]\to A\{R\}$ is $(p,d)$-completely faithfully flat, so is $A[R]\to A\{R\}_{\infty}$. Therefore, we obtain that $\mathbb{Z}_{(p)}[R]\to Z\{R\}$ is $p$-completely faithfully flat, and in particular, so is $R\to Z\{R\}\otimes_{Z_{(}p)[R]}R$. Since $R$ is $p$-torsion free, so is $Z\{R\}\otimes_{Z_{(p)}[R]}R$. By [MSTWW]*Lemma A.2, we obtain that $T(R)$ is $p$-torsion free, as desired. ∎
Appendix B Test ideal of graded ring
In this section, we prove Theorem B.1, which is a technical result to use the proof of Theorem 6.3. The proof of Theorem B.1 is taught by Kenta Sato.
Theorem B.1.
Let $S:=\bigoplus_{m\in\mathbb{Z}_{\geq 0}}S_{m}$ be a Noetherian graded ring such that $S_{0}$ is a field of characteristic $p>0$. We assume $S$ is normal, quasi-Gorenstein and $a(S)<0$ and set $\mathfrak{m}:=S_{>0},d:=\dim{S}$. We further assume $\operatorname{Spec}{S}\backslash\{\mathfrak{m}\}$ is strongly $F$-regular. Then we have
$0^{*}_{H^{d}_{\mathfrak{m}}(S)}=\cap_{e\geq 0}\operatorname{Ker}(H^{d}_{\mathfrak{m}}(S)\xrightarrow{F^{e}}H^{d}_{\mathfrak{m}}(F^{e}_{*}S)).$
Proof.
Since $S$ is quasi-Gorenstein, we obtain $\omega_{S}\simeq S(\nu)$ as graded $S$-modules, then $\nu<0$ by $a(S)<0$. By the isomorphism, the trace map $F_{*}\omega_{S}^{(p)}\to\omega_{S}$ is corresponding to $F_{*}(S(\nu))^{(p)}\to S(\nu)$, which is denoted by $u$. By the Matlis duality, it is enough to show that $\tau(S)=\bigcap_{e\in\mathbb{Z}_{\geq 0}}\mathrm{Im}(u^{e}\colon F^{e}_{*}S\to S)$, where $\tau(S)$ is the test ideal of $S$. First, let us show the following claim.
Claim B.2.
We have $u^{e}(F^{e}_{*}S_{n})=0$ if $n\not\equiv\nu\ \textrm{mod}\ p^{e}$ for every positive integer $e$.
Proof.
Since $u^{e}\colon F^{e}_{*}(S(\nu)^{(p^{e})})\to S(\nu)$ is a morphism of graded $S$-modules, for an integer $m\geq 0$,
$(u^{e})^{-1}(S_{m})=(u^{e})^{-1}(S(\nu)_{-\nu+m})\subseteq F^{e}_{*}(S(\nu))^{(p^{e})}_{-\nu+m}=F^{e}_{*}S_{(1-p^{e})\nu+p^{e}m}.$
In particular, if $u^{e}(F^{e}_{*}a)\neq 0$ for a homogeneous element $a\in S$, then $\deg(a)\equiv\nu\ \textrm{mod}\ p^{e}$, as desired. $\blacksquare$
Next, let us show the following claim.
Claim B.3.
Let $\mathfrak{q}\subseteq S$ be a $\mathfrak{m}$-primary ideal and a rational number $01>t>$ 0. Then there exists a positive integer $e_{0}$, we have $u^{e}(F^{e}_{*}(\mathfrak{q}\cdot\mathfrak{m}^{\lceil t(p^{e}-1)\rceil}))=u^{e}(F^{e}_{*}S)$ for $e\geq e_{0}$.
Proof.
We may assume $\mathfrak{q}=\mathfrak{m}^{n}$ for some positive integer $n$. We take a rational number $01>t>$ 0. We take $e_{0}$ such that $n+t(p^{e_{0}}-1)\leq p^{e_{0}}+\nu-1$ and $e\geq e_{0}$. Since $\mathfrak{m}^{l}\subseteq S_{\geq l}$ for every positive integer $l$, we have
$u^{e}(F^{e}_{*}\mathfrak{m}^{n+\lceil t(p^{e}-1)\rceil})\subseteq u^{e}(F^{e}_{*}S_{\geq n+\lceil t(p^{e}-1)\rceil}).$
Thus, it is enough to show that $u^{e}(F^{e}_{*}S_{\leq n+\lceil t(p^{e}-1)\rceil})=0$. We have
$u^{e}(F^{e}_{*}S_{\leq n+\lceil t(p^{e}-1)\rceil})\overset{(\star_{1})}{\subseteq}u^{e}(F^{e}_{*}S_{\leq p^{e}+\nu-1})\overset{(\star_{2})}{=}0,$
where $(\star_{1})$ follows from the choice of $e_{0}$ and $(\star_{2})$ follows from Claim B.2, as desired. $\blacksquare$
Next, let us show the following claim:
Claim B.4.
For every rational number $01>t>$ 0, we have
$\sigma(S,\mathfrak{m}^{t})=\sigma(S)=u^{e}(F^{e}_{*}S),$
where $\sigma(S,\mathfrak{m}^{t})$ and $\sigma(S)$ are defined in [FST].
Proof.
By [FST]*Remark 14.5(1), we obtain $\sigma(S)=u^{e}(F^{e}_{*}S)$ for large enough $e$. Furthermore, for enough large $e$, we have
$\sigma(S)$ $\supseteq\sigma_{1}(S,\mathfrak{m}^{t})=\sum_{e^{\prime}\geq 1}u^{e^{\prime}}(F^{e^{\prime}}_{*}\overline{\mathfrak{m}^{\lceil t(p^{e^{\prime}}-1)\rceil}})$
$\supseteq u^{e}(F^{e}_{*}\mathfrak{m}^{\lceil t(p^{e}-1)\rceil})\overset{(\star_{3})}{=}u^{e}(F^{e}_{*}S),$
where $(\star_{3})$ follows from Claim B.3. Therefore, we have $\sigma_{1}(S,\mathfrak{m}^{t})=\sigma(S)$. Furthermore, since $S$ is $F$-pure outside of $\mathfrak{m}$, the ideal $\sigma(S)$ is $\mathfrak{m}$-primary. Thus, for large enough $e$, we have
$\sigma(S)$ $=\sigma_{2}(S)\supseteq\sigma_{2}(S,\mathfrak{m}^{t})=\sum_{e^{\prime}\geq 1}u^{e^{\prime}}(F^{e^{\prime}}_{*}\sigma_{1}(S,\mathfrak{m}^{t})\overline{\mathfrak{m}^{\lceil t(p^{e^{\prime}}-1)\rceil}})$
$\supseteq u^{e}(F^{e}_{*}\sigma_{1}(S,\mathfrak{m}^{t})\mathfrak{m}^{\lceil t(p^{e}-1)\rceil})\overset{(\star_{4})}{=}u^{e}(F^{e}_{*}S)=\sigma(S),$
where $(\star_{4})$ follows from Claim B.3, thus we have $\sigma_{2}(S,\mathfrak{m}^{t})=\sigma_{1}(S,\mathfrak{m}^{t})=\sigma(S)$. Therefore, we have $\sigma(S,\mathfrak{m}^{t})=\sigma(S)$, as desired. $\blacksquare$
Finally, let us show the assertion. We take rational numbers $01>t>\varepsilon>$ 0, then we have
$\tau(S)\subseteq\sigma(S)\overset{(\star_{5})}{=}\sigma(S,\mathfrak{m}^{t})\overset{(\star_{6})}{\subseteq}\tau(S,\mathfrak{m}^{t-\varepsilon})\subseteq\tau(S),$
where $(\star_{5})$ follows from Claim B.4 and $(\star_{6})$ follows from [FST]*Proposition 14.10(4) and the fact that $S$ is strongly $F$-regular outside of $\mathfrak{m}$. Therefore, we obtain that
$\tau(S)=\sigma(S)=\bigcap_{e\in\mathbb{Z}_{\geq 0}}\mathrm{Im}(u^{e}\colon F^{e}_{*}S\to S),$
as desired. ∎
Appendix C Proof of Fedder type criterion
In this section, we provide a proof of Theorem 4.12. The proof of Theorem 4.12 is almost identical to a proof of [TWY]*Theorem A, however, for the convenience of the reader, we provide a proof.
Notation C.1.
Let $(A,\mathfrak{m})$ be a regular local ring with finite ring homomorphism $\phi\colon A\to A$ such that $\phi$ is a lift of Frobenius. We assume $p\in\mathfrak{m}\backslash\mathfrak{m}^{2}$ and set $\overline{A}:=A/pA$, then $\overline{A}$ is regular. We fix a generator $u\in\operatorname{Hom}_{\overline{A}}(F_{*}\overline{A},\overline{A})$. Let $f_{1},\ldots,f_{r}$ be a regular sequence in $A$ and we set $I:=(f_{1},\ldots,f_{r})\overline{A}$ and $f:=f_{1}\cdots f_{r}$. We assume $R:=A/(f_{1},\ldots,f_{r})$ is $p$-torsion free and set $\overline{R}:=R/pR$. By Theorem 3.7, we obtain the isomorphism
(C.1) $\operatorname{Hom}_{A}(Q_{A,n},\overline{A})\simeq\operatorname{Hom}_{A}(F^{n}_{*}\overline{A},\overline{A})\oplus\cdots\oplus\operatorname{Hom}_{A}(F_{*}\overline{A},\overline{A})\overset{(\star_{1})}{\simeq}F_{*}^{n}\overline{A}\oplus\cdots\oplus F_{*}\overline{A},$
where $(\star_{1})$ is given by
$F^{e}_{*}\overline{A}\xrightarrow{}\operatorname{Hom}_{A}(F^{e}_{*}\overline{A},\overline{A})\ ;\ F^{e}_{*}a\mapsto(F^{e}_{*}b\mapsto u^{e}(F^{e}_{*}(ab)))$
for each $n\geq e\geq 1$. The homomorphism corresponding to $(F^{n}_{*}g_{1},\ldots,F_{*}g_{n})$ is denoted by $\varphi_{(g_{1},\ldots,g_{n})}$.
Lemma C.2.
We use Notation C.1. Let $n$ be a positive integer. Then $R$ is $n$-quasi-$F$-split if and only if there exists an $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that
  1. (a)
    $\psi(1)$ is a unit and
  2. (b)
    $\psi(\operatorname{Ker}(Q_{A,n}\to Q_{R,n}))\subseteq I$.
Proof.
Since $\phi$ is finite and $A$ is regular, the module $\phi^{e}_{*}A$ is a free $A$-module for each positive integer $e$. By Theorem 3.7, the $A$-module $\phi_{*}W_{n}(A)$ is also free, and in particular, the $\overline{A}$-module $Q_{A,n}$ is free. First, we assume $R$ is $n$-quasi-$F$-split, then there exists $\overline{A}$-module homomorphism $\overline{\psi}\colon Q_{R,n}\to\overline{R}$ such that $\overline{\psi}(1)=1$. Since $Q_{A,n}$ is a free $\overline{A}$-module, there exists $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that the diagram
[…]
and in particular, $\psi$ satisfies the conditions (a) and (b). Next, we assume that there exists an $\overline{A}$-module homomorphism $\psi\colon Q_{A,n}\to\overline{A}$ such that $\psi$ satisfies the conditions (a) and (b). Thus, the homomorphism $\psi$ induces the $\overline{R}$-module homomorphism $\overline{\psi}\colon Q_{R,n}\to\overline{R}$ such that $\overline{\psi}(1)$ is a unit. Replacing $\overline{\psi}$ by $\overline{\psi}(1)^{-1}\cdot\overline{\psi}$, we obtain a splitting of $\Phi_{R,n}\colon\overline{R}\to Q_{R,n}$, as desired. ∎
Lemma C.3.
We use Notation C.1. Let $n$ be an integer with $n\geq 2$. We take $\varphi\in\operatorname{Hom}_{\overline{A}}(Q_{A,n},\overline{A})$ and $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$. By the isomorphism $Q_{A,n}\simeq F_{*}Q_{A,n-1}\oplus F_{*}\overline{A}$ in the proof of Theorem 3.7, we obtain the homomorphism $\psi\in\operatorname{Hom}_{\overline{A}}(F_{*}Q_{A,n-1},\overline{A})$ and $h\in\overline{A}$ such that
$\varphi(\alpha)=\psi(F_{*}\Delta_{W_{n}}(\alpha))+u(F_{*}(ha_{0}))$
for every $\alpha=(a_{0},\ldots,a_{n-1})\in Q_{A,n}$. Furthermore, we assume $\varphi$ satisfies the condition (b) in Lemma C.2.
  1. (1)
    There exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g_{s}\equiv f^{p^{n+1-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n+1-s}]}$ for every $n\geq s\geq 1$,
  2. (2)
    We use $h_{1},\ldots,h_{n}$ in (1). If $n=2$ or $p\neq 2$, then we have
    (C.2) $h_{n}-u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))\in f^{p-1}\overline{A}+I^{[p]},$
    and if $n\geq 3$ and $p=2$, then we have
    (C.3) $h_{n}-u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))-u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in f^{p-1}\overline{A}+I^{[p]}.$
Proof.
Let us show (1) by induction on $n\geq 2$. We note that $h_{n}=g_{n}$ satisfies the desired condition for $s=n$. Since $\psi$ coincides with the decomposition
[…]
and $\varphi$ satisfies the condition (b) in Lemma C.2, the homomorphism $\psi$ induces the $\overline{R}$-module homomorphism $\psi_{R}\colon F_{*}Q_{R,n-1}\to\overline{R}$. Since $\overline{R}$ is Gorenstein, the homomorphism $\psi_{R}$ factors through the homomorphism induced by $u(F_{*}(f^{p-1}\cdot-))\colon F_{*}\overline{A}\to\overline{A}$. Therefore, there exists $\varphi^{\prime}\in\operatorname{Hom}_{\overline{A}}(Q_{A,n-1},\overline{A})$ such that $\psi$ and $u(F_{*}(f^{p-1}\cdot-))\circ\varphi^{\prime}$ induce same homomorphism $Q_{R,n}\to\overline{R}$ and $\varphi^{\prime}$ satisfies the condition (b) in Lemma C.2. We take $g^{\prime}_{1},\ldots,g^{\prime}_{n-1}$ such that $\varphi^{\prime}=\varphi_{(g^{\prime}_{1},\ldots,g^{\prime}_{n-1})}$. By the isomorphism
$\operatorname{Hom}_{\overline{A}}(F_{*}Q_{A,n-1},\overline{A})\simeq F^{n}_{*}\overline{A}\oplus\cdots\oplus F^{2}_{*}\overline{A},$
the homomorphism $\psi$ is corresponding to $(F^{n}_{*}g_{1},\ldots,F^{2}_{*}g_{n-1})$ and $u(F_{*}(f^{p-1}\cdot-))\circ\varphi^{\prime}$ is corresponding to $(F^{n}_{*}f^{p^{n-1}(p-1)}g^{\prime}_{1},\ldots,F^{2}_{*}f^{p(p-1)}g^{\prime}_{n-1})$. Since they induce the same homomorphism, we have $g_{s}\equiv f^{p^{n-s}(p-1)}g^{\prime}_{s}\ \textrm{mod}\ I^{[n+1-s]}$ for every $n-1\geq s\geq 1$. For $n=2$, we have $g_{1}\equiv f^{p^{2}-p}g^{\prime}_{1}\ \textrm{mod}\ I^{[p^{2}]}$, thus $h_{1}=g^{\prime}_{1}$ satisfies the desired condition. For $n\geq 3$, by induction hypothesis, we have
$g_{s}\equiv f^{p^{n-s}(p-1)}g^{\prime}_{s}\equiv f^{p^{n-s}(p-1)}f^{p^{n-s}-p}h_{s}\ \textrm{mod}\ f^{p^{n-s}(p-1)}I^{[p^{n-s}]}\subseteq I^{[p^{n+1-s}]}$
for $n-1\geq s\geq 1$, as desired.
Next, let us show assertion (2). We take $r\geq j\geq 1$ and $a\in A$. Since $\varphi$ satisfies the condition (b), we obtain
(C.4) $\varphi([af_{j}])=\sum_{s=1}^{n}u^{n+1-s}(F^{n+1-s}_{*}(g_{s}\Delta_{n-s}(af_{j})))\in I.$
where we define $\Delta_{0}(af_{j})=af_{j}$. For $n-s\geq 3$ or $p\neq 2$, we have $g_{s}\Delta_{n-s}(af_{j})\in I^{[p^{n+1-s}]}$ by 3.6 and the assertion (1). Furthermore we have
$g_{n-1}\Delta_{1}(af_{j})$ $\equiv h_{n-1}f^{p^{2}-p}\Delta_{1}(af_{j})\ \textrm{mod}\ I^{[p^{2}]}$
$=h_{n-1}\Delta(f^{p-1}af_{j})-h_{n-1}a^{p}f_{j}^{p}\Delta_{1}(f^{p-1})$
$\equiv-h_{n-1}a^{p}f_{j}^{p}\Delta_{1}(f^{p-1})\ \textrm{mod}\ I^{[p^{2}]}.$
Therefore, if $n=2$ or $p\neq 3$, we have
$u(F_{*}(af_{j}(g_{n}-u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1}))))))\in I$
for every $a\in\overline{A}$ and $r\geq j\geq 1$. Therefore, we have
$g_{n}-u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))\in(I^{[p]}\colon I)=f^{p-1}\overline{A}+I^{[p]}.$
On the other hand, if $n\geq 3$, $p=2$ and $s\geq 3$, then we have
$g_{n-s}\Delta_{s}(af_{j})\equiv f^{p^{s+1}-p}h_{n-s}(af_{j})^{p^{s}-p^{2}}\Delta_{1}(af_{j})^{p}\equiv 0\ \textrm{mod}\ I^{[p^{s+1}]},$
and if $n\geq 3,p=s=2$, then we have
$g_{n-2}\Delta_{2}(af_{j})$ $\equiv f^{p^{3}-p}h_{n-2}\Delta_{1}(af_{j})^{p}\ \textrm{mod}\ I^{[p^{3}]}$
$\equiv h_{n-2}f^{p-1}(f^{p^{2}-p}\Delta_{1}(af_{j}))^{p}\ \textrm{mod}\ I^{[p^{3}]}$
$\equiv h_{n-2}f^{p-1}a^{p^{2}}f_{j}^{p^{2}}\Delta_{1}(f^{p-1})^{p}\ \textrm{mod}\ I^{[p^{2}]}.$
Therefore, by the same argument in the case of $n=2$ or $p\neq 3$, we have
$h_{n}-u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))-u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in f^{p-1}\overline{A}+I^{[p]},$
as desired. ∎
Proof of Theorem 4.12.
First, we note that sequences $\{I_{n}\}$ and $\{I^{\prime}_{n}\}$ are increasing. Indeed, $I_{1}\subseteq I_{2}$ by definition and if $I_{n-1}\subseteq I_{n}$, then
$I_{n}=u(F_{*}(I_{n-1}\Delta_{1}(f^{p-1})))+I_{1}\subseteq u(F_{*}(I_{n}\Delta_{1}(f^{p-1})))+I_{1}=I_{n+1}$
for every $n\geq 2$. Furthermore, $I^{\prime}_{n}\subseteq I^{\prime}_{n+1}$ for every $n\geq 1$ follows from the same argument. Let us show assertion (1). First, we consider the following claim.
Claim C.4.
Let $g\in A$ and $n\geq 1$. Then $g\in I_{n}$ if and only if there exists $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2 and $g_{n}\equiv g\ \textrm{mod}\ I_{n-1}$, where if $n=1$, we regard $I_{0}$ as $(0)$.
Proof.
Let us show assertion by induction on $n$. For $n=1$, since $I_{1}=(I^{[p]}\colon I)$, $g\in I_{1}$ if and only if $u(F_{*}(g\cdot-))$ satisfies the condition (b) in Lemma C.2, as desired. Next, we assume $n\geq 2$. We take $g\in I_{n}$, then there exists $g^{\prime}\in I_{n-1}$ such that $g-u(F_{*}(g^{\prime}\Delta_{1}(f^{p-1})))\in I_{1}$. By the induction hypothesis, there exist $g^{\prime}_{1},\ldots,g^{\prime}_{n-1}$ such that $\varphi_{(g_{1}^{\prime},\ldots,g^{\prime}_{n-1})}$ satisfies the condition (b) in Lemma C.2, $g^{\prime}_{n-1}\equiv g^{\prime}\ \textrm{mod}\ I_{n-2}$, then by Lemma C.3, there exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g^{\prime}_{s}\equiv f^{p^{n-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n-s}]}$ for each $n-1\geq s\geq 1$. We define $\varphi^{\prime}$ by
$\varphi^{\prime}\colon F_{*}Q_{A,n-1}\xrightarrow{F_{*}\varphi_{(g_{1}^{\prime},\ldots,g^{\prime}_{n-1})}}F_{*}\overline{A}\xrightarrow{u(F_{*}f^{p-1}\cdot-)}\overline{A}$
and $\varphi$ by
$\varphi\colon Q_{A,n}\xrightarrow{\sim}F_{*}Q_{A,n-1}\oplus F_{*}\overline{A}\xrightarrow{(\varphi^{\prime},u(F_{*}g_{n}\cdot-))}\overline{A},$
where
$g_{n}:=\begin{cases}u(F_{*}(g^{\prime}_{n-1}\Delta_{1}(f^{p-1})))&\textup{$n=2$ or $p\neq 2$}\\ u(F_{*}(g^{\prime}_{n-1}\Delta_{1}(f^{p-1})))+u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))&\textup{otherwise}.\end{cases}$
Since we have
$u(F_{*}(g^{\prime}_{n-1}\Delta_{1}(f)))\equiv u(F_{*}(g^{\prime}\Delta_{1}(f)))\equiv g\ \textrm{mod}\ I_{n}$
and $u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in I_{2}$, we obtain $g_{n}\equiv g\ \textrm{mod}\ I_{n}$. Thus, it is enough to show that $\varphi$ satisfies the condition (b) in Lemma C.2. We take $g_{1},\ldots,g_{n-1}$ such that $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$, then we have
$g_{s}=f^{p^{n-s}(p-1)}g^{\prime}_{s}\equiv f^{p^{n+1-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n+1-s}]}$
for $n-1\geq s\geq 1$. By the proof of Lemma C.3, for $a\in\overline{A}$, $r\geq j\geq 1$, we have
$\begin{array}[]{ll}\varphi([af_{j}])\equiv u(F_{*}(af_{j}(g_{n}-u(F_{*}(g^{\prime}_{n-1})))))\ \textrm{mod}\ I&\textup{ if $n=2$ or $p\neq 2$}\end{array}$
and
$\varphi([af_{j}])\equiv u(F_{*}(af_{j}(g_{n}-u(F_{*}(g^{\prime}_{n-1}\Delta_{1}(f^{p-1})))-u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2}))))))\ \textrm{mod}\ I$
otherwise. By the choice of $g_{n}$, we have $\varphi([af_{j}])\in I$ for all $a\in\overline{A}$ and $r\geq j\geq 1$. Furthermore, since we have
$[\sum_{j=1}^{r}a_{j}f_{j}]\equiv\sum_{j=1}^{r}[a_{j}f_{j}]\ \textrm{mod}\ V(W_{n-1}((f_{1},\ldots,f_{r})))$
and
$\varphi(V\alpha)=\varphi^{\prime}(\alpha)\in I$
for $\alpha\in W_{n-1}((f_{1},\ldots,f_{r}))$, we obtain that $\varphi$ satisfies the condition (b) in Lemma C.2, as desired.
Next, let us show the converse direction, thus we take $g_{1},\ldots,g_{n}\in\overline{A}$ such that the corresponding homomorphism $\varphi:=\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2, then it is enough to show that $g_{n}\in I_{n}$. For $n=1$, we have $g_{1}\in(I^{[p]}\colon I))=I_{1}$, as desired, thus we assume $n\geq 2$. By Lemma C.3, there exist $h_{1},\ldots,h_{n}\in\overline{A}$ such that $g_{s}\equiv f^{p^{n+1-s}-p}h_{s}\ \textrm{mod}\ I^{[p^{n+1-s}]}$, and in particular, $h_{n}\equiv g_{n}\ \textrm{mod}\ I^{[p]}$. By the proof of Lemma C.3 and the induction hypothesis, we obtain that $h_{n-1}\in I_{n-1}$, thus we have $u(F_{*}(h_{n-1}\Delta_{1}(f^{p-1})))\in I_{n}$. Furthermore, by the construction of $I_{2}$, we have
$u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in I_{2}.$
Therefore, by Lemma C.3 (2), we have $g_{n}\in I_{n}$, as desired. $\blacksquare$
Finally, let us show the assertion. We have
$\mathrm{Im}(\operatorname{Hom}_{\overline{R}}(Q_{R,n},\overline{R})\xrightarrow{\mathrm{ev}}\overline{R})$
[…]
$\overset{(\star_{2})}{\equiv}$ $u(F_{*}I_{n})\cdot\overline{R}\ \textrm{mod}\ u(F_{*}I_{n-1})\cdot\overline{R},$
where $(\star_{1})$ follows from the proof of Lemma C.2 and $(\star_{2})$ follows from Claim C.4 and $\varphi_{(g_{1},\ldots,g_{n})}(F_{*}1)=u(F_{*}g_{n})$. In conclusion, we have $\operatorname{ht}(R):=\inf\{n\geq 1\mid I_{n}\nsubseteq\mathfrak{m}^{[p]}\}$, as desired.
Next, let us show assertion (2). We define the sequence of ideals $\{I^{e}_{n}\}$ by $I^{e}_{1}:=f^{p-1}u^{e}(F^{e}_{*}f^{p^{e}-1}\overline{A})+I^{[p]}$ and
$I^{e}_{n}:=u(F_{*}(I^{e}_{n-1}\Delta_{1}(f^{p-1}))+I_{1},$
inductively. Then $\{I^{e}_{n}\}$ is increasing and $I_{n}\subseteq I^{e}_{n+1}$ for every $n,e\geq 1$. First, we consider the following claim.
Claim C.5.
Let $g\in A$ and $n,e\geq 1$. Then $g\in I^{e}_{n}$ if and only if there exists $g_{1},\ldots,g_{n}\in\overline{A}$ such that $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2, there exists $h\in f^{p^{e+n}-1}\overline{A}$ with $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, and $g_{n}\equiv g\ \textrm{mod}\ I^{e}_{n-1}$, where if $n=1$, we regard $I^{e}_{0}$ as $(0)$.
Proof.
For $n=1$, it follows from definition and the fact that the condition (b) in Lemma C.2 is equivalent to $g_{1}\in f^{p-1}\overline{A}+I^{[p]}$. We assume $n\geq 2$. We take $g\in I^{e}_{n}$, then there exists $g^{\prime}\in I^{e}_{n-1}$ such that $g-u(F_{*}(g^{\prime}\Delta_{1}(f^{p-1})))\in I_{1}$. By induction hypothesis, there exist $g_{1}^{\prime},\ldots,g^{\prime}_{n-1}$ such that $\varphi_{(g^{\prime}_{1},\ldots,g^{\prime}_{n-1})}$ satisfies the condition (b) in Lemma C.2, we have $u^{e}(F^{e}_{*}h^{\prime})=g^{\prime}_{1}$ for some $h^{\prime}\in f^{p^{e+n-1}-1}\overline{A}$, and $g^{\prime}_{n-1}\equiv g^{\prime}\ \textrm{mod}\ I^{e}_{n-2}$. By the proof of Claim C.4, if we put $g_{s}:=f^{p^{n-s}(p-1)}g^{\prime}_{s}$ for every $n-1\geq s\geq 1$ and
$g_{n}:=\begin{cases}u(F_{*}(g^{\prime}_{n-1}\Delta_{1}(f^{p-1})))&\textup{$n=2$ or $p\neq 2$}\\ u(F_{*}(g^{\prime}_{n-1}\Delta_{1}(f^{p-1})))+u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))&\textup{otherwise},\end{cases}$
then $\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) Lemma C.2 and $g\equiv g_{n}\ \textrm{mod}\ I^{\prime}_{n}$. We note that $I_{2}\subseteq I^{e}_{n}$ if $n\geq 3$. Furthermore, we have
$u^{e}(F^{e}_{*}(f^{p^{e+n}-p^{e+n-1}}h^{\prime}))=f^{p^{n}-p^{n-1}}u^{e}(F^{e}_{*}h^{\prime})=f^{p^{n}-p^{n-1}}g^{\prime}_{1}=g_{1},$
as desired.
Next, let us show the converse direction, thus we take $g_{1},\ldots,g_{n}\in\overline{A}$ such that the corresponding homomorphism $\varphi:=\varphi_{(g_{1},\ldots,g_{n})}$ satisfies the condition (b) in Lemma C.2 and there exists $h\in f^{p^{e+n}-1}\overline{A}$ such that $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, then it is enough to show that $g_{n}\in I^{e}_{n}$. For $n=1$, it follows from the definition of $I^{e}_{1}$. For $n\geq 2$, we take $h_{1},\ldots,h_{n}$ as in Lemma C.3, then the assertion follows from $h_{n-1}\in I^{e}_{n-1}$ by the induction hypothesis and
$u(F_{*}(f^{p-1}\Delta_{1}(f^{p-1})u(F_{*}h_{n-2})))\in I_{2}\subseteq I^{e}_{n}$
for $n\geq 3$. $\blacksquare$
If $\operatorname{ht}(R)=\infty$, then $I_{n}^{\prime}\subseteq I_{n}\subseteq\mathfrak{m}^{[p]}$ by (1) for every positive integer $n$, thus we may assume $n:=\operatorname{ht}(R)<\infty$. We note that $I^{\prime}=u^{e}(F^{e}_{*}f^{p^{e}-1}\overline{A})$ for some $e\geq 1$ by [Gabber]*Lemma 13.1, thus if we take a such $e$, then we have $I^{\prime}_{n}=I^{e}_{n}$. First, we assume $R$ is quasi-$(F,F^{\infty})$-split. Then there exist $\varphi\colon Q_{A,n}\to\overline{A}$ satisfying conditions (a), (b) in Lemma C.2 and $h\in f^{p^{e+n}-1}\overline{A}$ such that we have
$u^{e+n}(F^{e+n}_{*}(ha^{p^{e}}))=\varphi(0,\ldots,0,a)$
for every $a\in\overline{A}$. If $\varphi$ corresponds to $(F^{n}_{*}g_{1},\ldots,F_{*}g_{n})$, then we have
$u^{e+n}(F^{e+n}_{*}(ha^{p^{e}}))=u^{n}(F^{n}_{*}(g_{1}a))$
for every $a\in\overline{A}$, and in particular, we have $u^{e}(F^{e}_{*}h)=g_{1}$. By Claim C.5, we have $g_{n}\in I^{e}_{n}=I^{\prime}_{n}$. Since $\varphi(1)=u(F_{*}g_{n})\notin\mathfrak{m}$, we have $I^{\prime}_{n}\nsubseteq\mathfrak{m}^{[p]}$, as desired. On the other hand, if $I^{e}_{n}=I^{\prime}_{n}\nsubseteq\mathfrak{m}^{[p]}$, then there exists $\varphi=\varphi_{(g_{1},\ldots,g_{n})}$ such that $\varphi$ satisfies the condition (b) in Lemma C.2, there exists $h\in f^{p^{e+n}-1}\overline{A}$ such that $u^{e}(F^{e}_{*}h)\equiv g_{1}\ \textrm{mod}\ I^{[p]}$, and $u(F_{*}g_{n})\notin\mathfrak{m}$ by Claim C.5. Then $(\varphi,u^{e+n}(F^{e+n}_{*}h\cdot-))\colon Q_{A,n}\oplus F^{e+n}_{*}\overline{A}\to\overline{A}$ induces the homomorphism $Q_{R,n,e}\to\overline{R}$ and $(\varphi,u(F^{e}_{*}h\cdot-))(1)$ is a unit, as desired. ∎